### v1

# Pushing JWST to the extremes: search and scrutiny of bright galaxy candidates at z15-30

Abstract
Aims. We investigate the galaxy UV Luminosity Function at to constrain early galaxy formation scenarios aimed at explaining the mild evolution of the UV LF bright-end found by JWST at z10-15.
Methods. We designed customized Lyman-break color selection techniques to identify galaxy candidates in the redshift ranges and . The selection was performed on the ASTRODEEP-JWST multi-band catalogs of the CEERS, Abell-2744, JADES, NGDEEP, and PRIMER survey fields, covering a total area of sq. deg.
Results. We identify nine candidates at , while no objects are found based on the color selection criteria. Despite exhibiting a 1.5 mag break, all the objects display multimodal redshift probability distributions across different SED-fitting codes and methodologies. The alternative solutions correspond to poorly understood populations of low-mass quiescent or dusty galaxies at z3-7. This conclusion is supported by the analysis of a NIRSpec spectrum recently acquired by the CAPERS program for one interloper object, which is confirmed to be a dusty ( mag) starburst galaxy at . We measured the UV luminosity function under different assumptions on the contamination level within our sample. We find that if even a fraction of the candidates is indeed at , the resulting UV LF points to a very mild evolution compared to estimates at , implying a significant tension with existing theoretical models. In particular, confirming our bright () candidates would require substantial revisions to the theoretical framework. In turn, if all these candidates will be confirmed to be interlopers, we conclude that future surveys may need ten times wider areas to select galaxies at . Observations in the F150W and F200W filters at depths comparable to those in the NIRCam LW bands are also required to mitigate contamination from rare red objects at z8.
1 Introduction
Since the beginning of its operations, JWST has easily enabled the detection of galaxies beyond , breaking the redshift barrier that was the consequence of the limited infrared (IR) sensitivity of HST, Spitzer and of ground–based telescopes. Several bright galaxies, more than expected on the basis of the observed evolution at or from theoretical models, were quickly detected by the very first Early Release Science observations (e.g., Naidu et al. 2022a; Castellano et al. 2022; Finkelstein et al. 2022). This result has been later statistically corroborated by wider-area surveys (e.g., Harikane et al. 2023; Castellano et al. 2023; McLeod et al. 2024; Donnan et al. 2024) and on the basis of spectroscopically confirmed samples (Harikane et al. 2024; Napolitano et al. 2025). In fact, spectroscopic confirmation and characterization have been extremely efficient at z10-12 (e.g., Arrabal Haro et al. 2023a, b; Roberts-Borsani et al. 2024b; Castellano et al. 2024; Napolitano et al. 2024) and up to (Carniani et al. 2024b, a), the current record holder.
This situation is in stark contrast with the poor constraints available at . Despite the fact that UV rest frame emission is, in principle, within NIRCam spectral coverage up to z30, attempts to identify galaxies at in existing surveys have led to very few candidates (Yan et al. 2023a, b; Leung et al. 2023; Austin et al. 2023; Conselice et al. 2024; Robertson et al. 2024; Kokorev et al. 2024; Whitler et al. 2025; Gandolfi et al. 2025; Pérez-González et al. 2025), none of which has yet been confirmed spectroscopically. The most striking example is object CEERS-93316, which appeared to be a strong candidate based on CEERS imaging data (Donnan et al. 2023; Harikane et al. 2023), but which eventually proved to be a red galaxy at whose photometry appears consistent with a Lyman-break galaxy due to a very unfortunate combination of a red continuum with extremely strong rest-optical line emission (Arrabal Haro et al. 2023b).
From an observational point of view, several effects conspire to make selection at 15 more difficult. On the one hand, objects become fainter and are detected in a smaller number of photometric bands, making the detection of spectral breaks and the constraints on the UV continuum less significant. On the other hand, the contamination in photometric samples is expected to increase with redshift, as true sources become rarer relative to potential contaminants (Vulcani et al. 2017), and this trend may be worsened by new classes of poorly characterized low/intermediate-redshift objects entering the selection criteria (Zavala et al. 2023; Pérez-González et al. 2023a, 2024c; Glazebrook et al. 2023; Bisigello et al. 2023, 2025; Rodighiero et al. 2023; Gandolfi et al. 2025).
Breaking the z15 barrier is fundamental for testing theoretical models of galaxy evolution and for approaching the epoch of formation of the first stars and first black holes. In fact, the different explanations that have been invoked to explain the mild evolution of the UV LF at z10 differ in their predictions at earliest times (Kokorev et al. 2024; Pérez-González et al. 2025). For instance, sustained high luminosity density beyond is favoured by changes in the initial mass function (IMF) or by an increased star-formation efficiency (e.g. Dekel et al. 2023; Trinca et al. 2024; Hutter et al. 2025; Mauerhofer et al. 2025), as well as by a rapid assembly of baryons (McGaugh et al. 2024), while the prediction of an earlier phase of dusty-enshrouded star-formation (e.g. Ferrara et al. 2023; Ziparo et al. 2023) or alternative dark energy or dark matter scenarios (Menci et al. 2024; Gandolfi et al. 2022) result in a sharp decline with redshift of the luminosity density.
In this paper we analyse the ASTRODEEP-JWST photometric sample (Merlin et al. 2024, M24 hereafter), which provides consistent measurements on the major JWST deep surveys, to select bright galaxy candidates at z15-30. We briefly present the dataset in Sect. 2. In Sect. 3 we describe our specific renditions of the Lyman-break technique and the results of our search for galaxy samples at and . We investigate in detail alternative low-redshift solutions of our candidates, and the spectroscopic properties of one newly confirmed interloper at z=6.56 in Sect. 4. The implications on the evolution of the UV LF and a comparison with theoretical models are presented in Sect. 5, while Sect. 6 explores the lessons learned from our analysis for designing future observations of galaxies at z15. The results are summarised in Sect. 7.
2 Observations and data analysis
We used JWST and HST photometric measurements from the ASTRODEEP-JWST catalogs presented in M24. We analysed the seven surveys comprising the public catalog release111http://www.astrodeep.eu/astrodeep-jwst-catalogs/: CEERS (ERS 1345, P.I. Finkelstein, Finkelstein et al. 2025); the JADES-GS (data release v2.0) and JADES-GN (v1.0) fields on the GOODS-South and GOODS-North footprints, respectively (GTO 1180 and GTO 1210, P.I. Eisenstein, Eisenstein et al. 2023) including FRESCO data (GO 1895, P.I. Oesch, Oesch et al. 2023); the first-epoch imaging of the NGDEEP field (Co-PIs Finkelstein, Papovich, Pirzkal, Bagley et al. 2024); the PRIMER (GO-1837, P.I. Dunlop) observations of the UDS and COSMOS fields in CANDELS (Grogin et al. 2011; Koekemoer et al. 2011); the A2744 field including JWST observations from GLASS-JWST (ERS 1324, P.I. Treu, Treu et al. 2022), UNCOVER (GO 2561, P.I. Labbé, Bezanson et al. 2022), DDT 2756 (P.I. Chen), and GO 3990 (P.I. Morishita, Morishita et al. 2024). The considered observations cover a significant range in both area and depth, from relatively wide surveys as PRIMER-UDS (250 sq. arcmin, 50% completeness at mag28.8 in the detection band), to deep pencil-beam pointings such as NGDEEP (9.5 sq. arcmin, mag30.8) and the lensed field A2744 (46 sq. arcmin, mag29.7). We refer to M24 for details on the survey properties and photometric techniques. Briefly, sources in all fields were detected with SExtractor (Bertin & Arnouts 1996) on a weighted average of the NIRCam F356W and F444W images, which is also used to measure total fluxes in Kron (1980) apertures using A-PHOT (Merlin et al. 2019). Fluxes in the other bands were measured by scaling the aforementioned total flux according to the colors measured within optimal apertures on PSF-matched images (see also Merlin et al. 2022; Paris et al. 2023). The seven fields comprise a total of 531173 objects in an area of 615 sq. arcmin, making the ASTRODEEP-JWST the largest publicly released JWST catalogs available to date. The available imaging datasets are slightly different in the various fields. In particular, JWST NIRCam F090W is missing, or was not public at the time of the catalog, in the CEERS, NGDEEP and over most of the A2744 area, while the medium band F410M filter is not available in NGDEEP and in the GLASS-JWST observations of A2744. Most importantly, as discussed in M24, the HST coverage is even less uniform both in terms of depth and number of available filters. For the present work we built the weighted average stacks of all the ACS and WFC3 HST bands available in each field, respectively. The SNR of all ASTRODEEP-JWST sources was measured within an aperture with a diameter of 2 times the PSF full-width at half maximum (FWHM) on both the HST stacks, and used as described in the following section to constrain non-detection blueward of the Lyman-break.
3 Lyman-break selection at z15-30
We describe here the approach that we have used to select high- candidates, that is a modified version of Lyman–break color selection criteria (Giavalisco 2002). The selected objects are then inspected through a full photometric redshift analysis to investigate their reliability.
3.1 Color selection criteria
We defined color selection criteria for galaxies in two different redshift ranges, z15-20 and z20-30 on the basis of mock catalogues of objects at z0-30 tailored to match the noise properties of our observations, as previously done for the z9-15 range by Castellano et al. (2022). Namely, we use two different simulations. The first simulation is based on a catalogue comprising objects at 0z5 over an area of 0.12 sq. deg. generated with the Empirical Galaxy Generator (EGG) code (Schreiber et al. 2017), which exploits empirical relations to reproduce the observed number counts and color distributions of galaxies at low and intermediate redshifts, including quiescent and dusty populations. The second simulation is based on the mock catalogs from the JAdes extraGalactic Ultradeep Artificial Realizations (JAGUAR, Williams et al. 2018), including predicted NIRCam fluxes for objects at 0.215 and stellar mass log(M/M⊙)6 over an area of 0.34 sq. deg. JAGUAR provides a complementary test with respect to EGG also thanks to the inclusion of emission lines in the predicted SEDs. We added sources at z5 (z15) to the EGG (JAGUAR) simulation following the evolving UV LF at z5-10 (Bouwens et al. 2021), assuming no evolution beyond and artificially boosting the number counts at z10 by a factor of 20 in order to provide sufficient statistics to design appropriate selection criteria. These high-redshift galaxies have been generated by randomly associating to each object a template from a library based on Bruzual & Charlot (2003, BC03 hereafter) models with metallicities 0.02 or 0.2 Z⊙, 0E(B-V)0.2 and a constant star-formation history (SFH) to predict the relevant photometry. The over-representation of high-redshift sources in the mock catalogs is taken into account by consistently scaling the relevant number counts when evaluating the selection criteria in terms of purity and completeness. Finally, we assessed the potential contamination by late-type dwarf stars using synthetic JWST photometry for the models by Marley et al. (2021) which include brown dwarfs and self-luminous extrasolar planets with and metallicity [M/H] from 0.5 to + 0.5. The brown dwarf models were normalized at 26.0F444W28.0 in 0.5 mag steps.
All the catalogues were perturbed by adding noise in order to reproduce the expected relation and scatter between magnitudes and errors in each band and in each of the analysed fields.
After extensive testing, we first define a detection threshold corresponding to SNR10 in the detection band used by M24, i.e. F356W+F444W. We also define the following selection criteria to identify objects at z15-20 minimizing contamination from low-redshift sources:
We require a signal-to-noise ratio SNR2.0 in the F090W (where available), F115W and F150W bands, and in both the ACS and WFC3 stack, with at most one of these bands blueward of the break having SNR1.5. In order to limit our sample to objects with continuous coverage redward of the Lyman break and to avoid spurious, single-band detections, we also require SNR2 in each of the F277W, F356W, F444W bands. All the adopted signal-to-noise ratios are measured in 2FWHM apertures.
Similarly, we find that objects at z20-28 are well identified by the following selection criteria, as shown in the bottom panel of Figure 1:
As above, we require SNR2 in the bands redward of the Lyman break (F356W, F444W), SNR2.0 blueward of the break (stacked ACS and WFC3 images, F090W, F115W, F150W and F200W bands), with at most one band blueward of the break having SNR1.5.
When analysing the observed dataset (Sect. 3.2), we will also exclude objects classified as spurious by M24, or after visual inspection, such as hot pixels and stellar spikes.
We find that the proposed diagrams efficiently select high-redshift targets up to z28, where the F356W-F444W color becomes 0.5 due to the Lyman-break entering the F356W band. Some contamination from low-redshift galaxies is evident from the redshift distribution of the selected objects (right panels in Fig. 1). We find a contamination rate of 0.05 objects/arcmin2 in the EGG-based simulations, 0.01 objects/arcmin2 in the JAGUAR-based simulation for the z15-20 selection, and 0.01 arcmin-2 in both simulations for the z20 selection. We do not find contamination by late-type dwarf stars in any of the proposed selection criteria, consistently with their expected colours and the detectable emission at 1m (Holwerda et al. 2018, 2024).
Our baseline selection criteria are meant to isolate a ?reference sample? that will be adopted to estimate the UV LF. In addition, we find that by lowering the color thresholds we can select additional very high-redshift targets, although with a significantly higher contamination rate (0.1/arcmin-2). An additional ?extended sample? of targets with (z15-20) and (z20) has been selected for potential follow-up and to characterize the properties of interloper populations in Sect. 4.2.
3.2 Selected sources over 0.2 deg2
Our baseline F200W dropout criteria for z15-20 yields a total of 12 objects. Among them, and consistently with previous works, we re-select as a z16 candidate the strong line-emitter object CEERS-93316 (ID=84213 in M24) with 4.9 (Arrabal Haro et al. 2023b). We also find that the only candidate selected in the A2744 field (ID=27713 in M24) is a transient source. In fact, object A2744_27713 is not detectable in the first epoch observations (June 2022) of the GLASS-JWST NIRCam parallel, but it is clearly detected in all LW bands from the second epoch dataset observed in November 2022. This object, whose redshift remains undetermined, masquerades as a F200W dropout because it falls in the chip gap of NIRCam SW second-epoch observations.
After excluding these two interlopers, we are left with an initial sample of 10 F200W dropouts. Of these, the only candidate selected in the PRIMER-UDS field (ID=56824 in M24) has been recently observed with NIRSpec PRISM observations by the CAPERS survey (GO-6368, P.I. M. Dickinson), yielding a spectroscopic redshift of , as discussed in more detail in Sect. 4.3.
The final ?reference? sample of F200W dropout candidates that we shall discuss in this paper is therefore made of 9 sources. We performed additional checks on their reliability. First of all, we measured the SNR within an aperture of 2 times the PSF FWHM on a stack of the NIRCam F090W, F115W and F150W bands available for each of them, finding that they are all non-detected at SNR2. We then computed the SNR on the NIRCam F090W, F115W and F150W bands in two other apertures measured by M24, namely an aperture with a radius of 0.1 arcsec (R01), and the one with a diameter of 3 times the PSF-FWHM. All objects are non-detected in these bands at SNR2 in all cases, except COSMOS_35731 having SNR=2.1 in the F150W band in the R01 aperture (compared to SNR=1.4 and SNR1 in the 2FWHM and 3FWHM apertures, respectively).
We estimated the MUV of the nine candidates by converting to the rest-frame the observed F277W magnitude assuming a redshift z=18 where our redshift selection function peaks, and their UV slope by fitting the F277W, F356W and F444W bands. We find three candidates brighter than M-21, with the brightest object in the sample COSMOS_107923 having M-22.7. Four of the objects have a red -2, the remaining ones being consistent with a flat or moderately blue UV slope (). We have measured their half-light radius by fitting the light distribution in the F277W band with GALIGHT (Ding et al. 2020; Birrer et al. 2021) assuming a Sersic (1968) profile with free index and fixing the redshift at the best-fit solution at . We find half-light radii consistent with those of observed in galaxies at z10-15 (Westcott et al. 2024; Ono et al. 2025). Our candidates have a typical half-light radius of 0.2-0.3 kpc, except COSMOS_31168, which is more extended (R0.5 kpc), and the compact source COSMOS_35731 (R1.5 pixel, amounting to 0.13 kpc) which is marginally resolved. Although its estimated size is comparable to that of other known sources at slightly lower redshifts (e.g., Ono et al. 2023; Tacchella et al. 2023) we verified both with GALIGHT and GALFIT (Peng et al. 2010) that a fit with a PSF-like profile provides comparable and residuals as the Sersic one.
We find a high field-to-field variance with 6 candidates found in PRIMER-COSMOS, 2 in CEERS, 1 in JADES-GN, and no sources selected in both JADES-GS and NGDEEP. As described above, neither A2744 nor PRIMER-UDS, which is the widest/shallowest among the considered fields, contribute to the final sample as the two selected sources are confirmed interlopers. Interestingly, candidates in both COSMOS and CEERS are close on the sky plane, such that if their high-redshift nature will be confirmed they would be at 1-5 physical Mpc distance from each other, possibly implying that they are part of distant overdensities.
We do not find any object meeting our reference F277W-dropout criteria.
In addition, we select extended samples of 26 F200W dropout objects, and of 19 F277W-dropout sources.
We show in Fig. 2 the position of all our candidates in the observed color-color diagrams. The IDs and main properties of objects in our reference F200W dropout sample are presented in Table 1, while their SED and NIRCam thumbnails are shown in Fig. 3. The SEDs and main properties of the objects in the extended samples are presented in the Appendix A.
We compared our samples with 15 samples selected by other groups in the fields analysed here. Our two F200W-dropout candidates in CEERS are not included in the sample of very red sources by Gandolfi et al. (2025), while 4 of their 5 objects potentially at z15 have indeed colours compatible with our selection window, but have not been included here because of SNR10 in the detection band, and in one case (their A-22691) a marginal (SNR2.2) detection in the ACS stacked image. The remaining source (U-53105) is not detected in the M24 catalog. Object NGD-z15a/NGDEEP-1369 presented as a z15.6 candidate by both Austin et al. (2023) and Leung et al. (2023) is matched to object ID=1301 in M24. It has colours consistent with our inclusive selection window but is not part of our ?extended sample? because of a detection at SNR2.5 in the stack of the ACS bands. Five of the sources selected by Pérez-González et al. (2025) in the MIDIS+NGDEEP observations, including the z19.6 candidate MDS025593 by Pérez-González et al. (2023b), are not covered by the first-epoch NGDEEP imaging used by M24, the remaining ones being non-detected except their MIDIS-z17-7 (ID=8412 in M24) which is at SNR4, i.e. well below our SNR=10 threshold. Finally, the 15 sources selected by Hainline et al. (2024) in the JADES fields with a counterpart in M24 do not fall within our colour selection window, and are detected at SNR10 except their JADES-GS-53.12692-27.79102 (ID=51718 in M24).
| ID | R.A. | Dec | F356W | M1500 | Re | zhigh | zlow | |
| deg. | deg. | AB | kpc | |||||
| CEERS_15937 | 214.944272 | 52.835847 | 27.15 0.09 | -21.14 | 0.263 0.046 | -1.59 0.35 | 17.2 | 4.6 |
| CEERS_17384 | 214.853243 | 52.773682 | 27.11 0.08 | -21.06 | 0.317 0.05 | -1.50 0.33 | 17.2 | 4.7 |
| JADES-GN_9538 | 189.191052 | 62.17421 | 28.90 0.13 | -19.49 | 0.227 0.043 | -1.63 0.40 | 16.1 | 2.8 |
| COSMOS_31168 | 150.180936 | 2.260756 | 27.67 0.16 | -20.70 | 0.519 0.124 | -2.05 0.39 | 16.2 | 4.0 |
| COSMOS_35731 | 150.133571 | 2.271020 | 28.54 0.09 | -20.05 | 0.134 0.025 | -1.92 0.17 | 15.9 | 2.8 |
| COSMOS_76919 | 150.184553 | 2.353510 | 28.71 0.18 | -19.74 | 0.289 0.065 | -2.32 0.44 | 16.1 | 4.0 |
| COSMOS_84213 | 150.167211 | 2.368995 | 28.14 0.18 | -20.35 | 0.199 0.053 | -2.03 0.40 | 15.6 | 4.0 |
| COSMOS_107923 | 150.107337 | 2.428780 | 25.48 0.10 | -22.68 | 0.219 0.024 | -1.22 0.26 | 16.3 | 3.3 |
| COSMOS_118438 | 150.196541 | 2.464192 | 27.86 0.19 | -20.61 | 0.256 0.077 | -2.27 0.47 | 17.1 | 5.0 |
a) ID, coordinates and F356W magnitudes from M24. The MUV and half-light radius () have been measured from the observed F277W band. The UV slope is obtained by fitting the F277W, F356W and F444W bands. The last two columns show the best-fit solutions obtained with zphot at () and ().
4 A closer scrutiny of the selected candidates
In this section we analyse in more detail the selected candidates to assess their reliability and evaluate the possibility that their peculiar colours are instead indicative of rare lower redshift interloper populations.
4.1 The photometric redshift probability distribution
Following a well-established practice, we have derived the photometric redshift of our candidates analysing their multi-wavelength photometry with a set of standard SED fitting tools. To alleviate the impact of specific flavours of the adopted techniques, and to broaden the range of spectral libraries explored, we have exploited four different codes: zphot (Fontana et al. 2000), EAzY (Brammer et al. 2008), BAGPIPES (Carnall et al. 2018, 2019) and CIGALE (Boquien et al. 2019). Rather than focusing on the best-fitting photometric redshift, we have used all these codes to compute the redshift probability distribution (). This distribution encapsulates the full information on the different potential solutions for an object with given photometry.
Briefly, we have run zphot adopting a set of synthetic models drawn from the BC03 library with a range of metallicities from to , , and a ?delayed? () star–formation history (see Santini et al. 2023, and M24). Nebular emission is self–consistently included following Schaerer & de Barros (2009) (see also Castellano et al. 2014), based on the template luminosity at ionizing frequencies. We have analysed our candidates with EAzY in two different ways as described in M24, i.e. 1) using only its standard set of semi-empirical templates, and 2) including also the recent set of templates from Larson et al. (2022) which are designed explicitly to reproduce the colors of high redshift galaxies. The Bagpipes runs exploit BPASS v. 2.2.1 stellar models with an upper-mass cutoff of the IMF of 300 M⊙ (Stanway & Eldridge 2018), and nebular emission computed self-consistently with CLOUDY (Ferland et al. 2013) as described by Carnall et al. (2018). Following Gandolfi et al. (2025), we increased the number of live points (i.e., the walkers used by Bagpipes in the Markov Chain Monte Carlo sampling) from the default 400 to 2000 to enhance sensitivity to strong line emitter solutions, and allowed the ionization parameter to reach log . We assume three different star-formation histories: delayed, double power law, and exponential SFH. Finally, we run CIGALE in two configurations. In the first case, we assume a SFH with a delayed component of age between 100 Myr and the age of the universe at each redshift, plus a constant burst of 10 Myr duration. The fraction of stellar mass formed in the recent burst is allowed to vary between 0 and 50% of the total assembled mass. We use BC03 templates including nebular emission, metallicity 0.02, 0.2, 1 Z⊙, and V-band extinction 05. The second CIGALE configuration exploits star-formation plus AGN templates where the stellar component is parametrized as described above and the AGN emission is based on the dale2014 module (Dale et al. 2014). The AGN fraction (fAGN), defined as the ratio of AGN luminosity to the total AGN and dust luminosities, is set as a free parameter. To summarise, from the combination of codes and assumptions adopted we have obtained eight different for all our candidates. The of the nine objects in our reference sample of F200W dropout candidates are shown in the upper panels of Fig. 3. We show in Fig. 4 the average for the extended samples of F200W- and F277W-dropout candidates.
Although there are specific differences among the various objects and the different codes, a number of general conclusions can be drawn from this analysis.
First, it is clear that all objects exhibit a solution, consistent with the color selection criteria adopted. The inferred redshifts tend to be reasonably similar among the codes, as the main spectral feature determining the high-redshift solution is the Lyman-break (coupled with the shape of the star–forming continuum immediately redward) that is essentially common in all recipes. However, all our candidates also show a lower-redshift solution, in all the runs analysed here. The low–redshift solutions are typically peaked at , suggesting that the strong observed break in the F200W band can also be ascribed to a break around the Balmer break/4000 rest-frame region, as we shall describe better in the following. In several cases z2 solutions are also viable, and generally preferred by BAGPIPES. Detailed inspection shows that this is generally due to a peculiar combination of strong emission lines that may conspire to reproduce the observed colours. We remark that, given the resulting , none of our candidates would pass a selection criterion such as 4 between different redshift solutions that has been often adopted to build luminosity functions at (e.g., Finkelstein et al. 2023; Harikane et al. 2023). Unfortunately, our candidates are simply too faint, and their photometry is built on a too small number of photometric bands with solid detections, to be unambiguously selected in the same manner. It is also clear that, while they overall provide a consistent picture of ?double peaked? solutions, there are significant differences between the adopted codes, both in terms of the breadth of the low-redshift solution and of the relative weight between the low- and high–redshift peaks. These differences arise from the different libraries adopted and probably from slight differences in the fitting procedure. This picture is confirmed by the average of the extended samples of F200W- and F277W-dropout candidates (Fig. 4).
We finally note that also depends crucially on the details of the adopted photometry. Clearly, but somewhat counterintuitively, the additional information provided by the HST photometry decreases the constraining power of . Because of the faintness of our candidates, in fact, the HST upper limits are easily satisfied by a wide class of solutions at low redshifts, eventually adding little contribution to the global while increasing the number of degrees of freedom and hence increasing the probability at low redshifts.
We take from these results three main lessons. The first is that all our objects are, in principle, credible candidates at , but none of them are solid enough to be unambiguously assigned these extreme redshifts. In addition, the differences among the various suggest that a detailed and sophisticated analysis built on their shape should be taken with caution, as they may depend on subtle details in the photometric measurements and on the spectral libraries adopted. Finally, as we will discuss below, the most important factor preventing a robust use of the in the selection process is our limited knowledge of the populations of faint, red interlopers.
4.2 The nature of potential interlopers
The photometric redshift distributions described above help us unveil the physical properties of the galaxies that may contaminate our selection criteria. Adopting for simplicity the results of the zphot code, we have inspected the physical properties corresponding to the models that populate the low redshift peaks in the . The best-fit SEDs at of the nine F200W dropout candidates are shown in Fig. 3.
In the case of the two candidates in CEERS, the low-redshift best fit solution is a ?quiescent? galaxy with low specific star–formation rate and large stellar age. The remaining sources, instead, are best reproduced by star–forming models with a very high specific star–formation rate and large dust attenuation, whose red continuum and strong emission lines yield colors compatible with our selection criteria.
These two kinds of solutions are representative of the general properties of galaxy templates that typically yield low-redshift solutions for our objects, as shown in Fig. 5. We compared the solutions that provide an acceptable fit with probability for the nine F200W dropout candidates to the locus of objects in the same redshift range from the JADES-GS field. The statistically acceptable models populate regions that have a small overlap with the ones occupied by the bulk of sources in the same redshift range. Consistently with the SEDs shown in Fig. 3, the templates cover a region in the versus plane that connects quiescent, low dust models (lower left corner) with highly star-forming, dusty (upper right corner) ones. Instead, the general distribution of galaxies in this plane shows that most of the objects tend to populate the region of intermediate and . The vs plane shows that acceptable solutions include z0-4 templates with a higher dust-extinction than ?typical? sources in the same redshift range, and, in particular, are consistent with potential contamination of the F200W dropout selection by very low-mass, dusty galaxies, as previously noted by Bisigello et al. (2023) and Gandolfi et al. (2025).
All these templates can reproduce the sharp break observed around in our objects. Because of the faintness of our candidates, the amplitude of the break (which is much larger in galaxies) cannot be properly measured with the existing photometry, leaving room for the ambiguity between the two redshift solutions.
We remark, however, that these potential low-redshift solutions correspond to objects that would be extremely interesting to investigate. These models have stellar masses in the range , sometimes even as low as . As long as quiescent galaxies are concerned, only objects with stellar mass above have been confirmed at (Carnall et al. 2024; Glazebrook et al. 2024; Weibel et al. 2024; Pérez-González et al. 2024b). The available estimates of the stellar mass function of quiescent galaxies do not extend to these low masses, especially at these redshifts (Santini et al. 2021). These sources may also host the low-mass SMBH that are fundamental to constrain early AGN-galaxy co-evolution (Pacucci et al. 2023).
Concerning dusty solutions, on the other hand, there is a general consensus that faint, low-mass galaxies at are essentially dust-free according to the strong correlations between dust-extinction and stellar mass (McLure et al. 2018; Bouwens et al. 2020). If (even some of) these candidates are instead low-mass, dusty galaxies as suggested by one class of solutions, they would correspond to a phase in galaxy evolution that has not been widely investigated so far. Recently, Bisigello et al. (2025) has spectroscopically confirmed a low mass, dusty galaxy at , suggesting that other similar objects (1-15% of the sources of similar mass and redshift) might be hidden in the photometric sample. Although rare, galaxies with rest-frame colors that approach our color selection criteria do exist in nature. As a result, we cannot exclude the possibility that some - or even all - our candidates are indeed intermediate redshift interlopers.
We cannot exclude that contamination from AGN is also present, although it seems less likely than from dusty or quiescent galaxies. In fact, when including AGN emission, CIGALE tends to increase rather than decrease the probability of solutions, in some cases indicating them as best-fit. Finally, we do not expect significant contamination from Little Red Dots (LRD) at z3-8, except possibly for the reddest objects with strong emission lines whose predicted colours are similar to those of dusty, high-sSFR galaxies (Killi et al. 2024; Pérez-González et al. 2024a).
4.3 The confirmed interloper UDS_56824 at =6.56
One of the ten objects originally selected, the F200W-dropout candidate object UDS_56824 at R.A.=34.454893 deg., Dec=-5.215586 deg (see Fig. 2), has been recently observed with the NIRSpec PRISM in the framework of the CAPERS program. The observation has been carried out by adopting a NRSIRS2 readout pattern, standard 3-shutter “slits”, and a 3-point nodding for a total integration time of 5690 secs. The data were reduced with the STScI Calibration Pipeline222https://jwst-pipeline.readthedocs.io/en/latest/index.html version 1.13.4 as described in detail in Arrabal Haro et al. (2023a) (see also Castellano et al. 2024; Napolitano et al. 2025). Although the spectrum covers only the region at 3m, the redshift can be accurately measured to be =6.560.01 from the weighted average centroid of the H, H, [O III] lines which are clearly detected (Fig. 6). We measured line fluxes with a Gaussian fit after linearly extrapolating the continuum emission at the line position (see Napolitano et al. 2024, 2025, for details) finding FHα=16.2 2.1 10, FHβ=2.1 0.9 10, F[OIII]4959=2.1 1.1 10, and F[OIII]5007=1.26 0.18 10.
The Balmer ratio implies a high dust attenuation with =0.850.46 when assuming a Calzetti et al. (2000) attenuation law. Its [O III]/H=6.02.7 puts the object in the star-forming region of the mass-excitation diagram (Juneau et al. 2014), although at the border with the AGN locus, such that a contribution from a dust-obscured active nucleus cannot be excluded.
We checked the ALMA archive finding that the position of UDS_56824 has been observed by project #2015.1.01074.S (PI H. Inami) that we analyse as follows. We start from the calibrated measurement set and use the CASA tclean function to create a continuum map using the four 2 GHz-wide spectral windows, with an effective central frequency of 343 GHz (i.e., 870 m). We apply natural weighting to the visibilities and test the effect of adding a uvtaper value to artificially increase the beam size. A 3 detection is found at the NIRCam position of UDS_56824, with an integrated flux density of mJy (see Fig. 6). This galaxy is fainter than all SCUBA-2-selected galaxies in UDS (Geach et al. 2017) and all previous ALMA detections in the field reported in Dudzevičiūtė et al. (2020), suggesting that a significant population of dusty galaxies at z6 may have been missed by previous submillimeter surveys.
The measured ALMA flux density was then combined with NIRCam photometry to perform an energy-balance SED fitting using BAGPIPES. During the fitting, the redshift was fixed to the spectroscopic value of z=6.56, while the dust extinction, AV, was allowed to vary freely between 1.8 and 5, as constrained by the Balmer decrement. The best-fit stellar mass and star formation rate are log = 9.8 0.18 and SFR= 40 M⊙ yr-1, respectively, implying that UDS_56824 lies above recent estimates of the main sequence at (Rinaldi et al. 2024; Cole et al. 2025). Additionally, we estimated the dust mass using standard relationships (e.g., Casey et al. 2019) and the following assumptions: , Td=25 K, and =1.8. This yields a dust mass of 5.5M⊙, close to 10% of the stellar mass - a remarkably high value given the redshift of the galaxy. Assuming a higher dust temperature of T K would lower the dust-to-stellar mass ratio to log -2.0, bringing it into better agreement with other high-redshift dusty galaxies (e.g., Ferrara et al. 2025b; Algera et al. 2024) and supporting a mild evolution toward higher dust temperatures at high redshifts (e.g., Mitsuhashi et al. 2024). Nevertheless, even in this scenario, this galaxy would stand out due to its high dust attenuation compared to UV-selected galaxies.
The combination of spectroscopic and photometric information demonstrates that UDS_56824 is an example of a highly attenuated high-redshift starburst galaxy. Its large UV attenuation 4.5 mag, makes it a significant outlier in the relation (McLure et al. 2018), similarly to CEERS-93316 (Arrabal Haro et al. 2023b) and CEERS-14821 (Bisigello et al. 2025). In fact, UDS_56824 shows that even at z6.5, some galaxies may be dominated by dust-obscured star formation (note that the uncorrected H-based SFR is 2-4 M⊙ yr-1). The other known interloper CEERS-93316 at z=4.9 is remarkably similar to UDS_56824 in terms of both extinction (0.6 mag) and SFR (30 M⊙ yr-1), but with a lower stellar mass log8.4-9 (Arrabal Haro et al. 2023b).
These findings are consistent with our analysis in Sect. 4.2 pointing toward dusty star-forming galaxies up to z7-8 as a potential source of contamination of the F200W-dropout sample. Similarly to CEERS-93316, the SED of UDS_56824 resembles that of a z15 LBG due to the combination of a red, attenuated continuum and line emission. However, it is apparent that a proper characterization of these interloper populations is extremely difficult and requires spectroscopic follow-up observations. In fact, a redshift of z6.5 is a secondary peak in our (Fig. 7) but the most likely alternative solution was a similarly dusty, star-forming object with log9 at z4 whose broad-band fluxes are boosted by a different combination of emission lines. Moreover, the properties of UDS_56824 appear extreme compared to the locus of alternative solutions of the other 15 candidates, while CEERS-93316 is more in line with the expectations (Fig. 5).
4.4 Critical assessment of the photometric sample
We exploit the analysis described above to perform a detailed evaluation of the objects in our reference sample of F200W-dropouts. We first note that four objects in our sample consistently show a probability for the high-redshift solution which is significantly lower than any of the low/intermediate redshift peaks: COSMOS_31168, COSMOS_76919, COSMOS_84213 and COSMOS_118438. In the case of COSMOS_31168 and COSMOS_84213, this is likely due to a marginal detection (SNR2) in one HST band each, albeit both objects are non-detected in both the ACS and WFC3 stacks that we used for our selection. In addition, both these sources as well as COSMOS_76919 show a drop in the F410M band similarly to UDS_56824, thus making dusty star-forming solutions more likely. On the contrary, COSMOS_118438 has a slightly higher flux in F410M than in F356W and F444W, which also leads SED-fitting codes to prefer strong-line emitting templates at z6. In fact, we tested that the probability of the solution increases after removing the F410M from the fit for all the aforementioned sources. The remaining five sources in our sample have high-redshift solutions as significant as the low-redshift ones in most of the SED-fitting runs. In particular, CEERS_15937, CEERS_17384 and COSMOS_107923 show a high consistency among all codes and recipes. However, there are compelling reasons to consider their reliability with caution. The two CEERS objects appear somewhat similar to CEERS-93316, as they have a clear peak at and a red continuum in the F277W/F356W/F444W bands, leading us to suspect that they are members of the same 5 overdensity (Naidu et al. 2022b; Arrabal Haro et al. 2023b) together with other known quiescent (de Graaff et al. 2025) and dusty (Zavala et al. 2023; Bisigello et al. 2025) galaxies. Finally, COSMOS_107923, albeit being the brightest in the sample falls in a region of relatively shallow image depth and therefore has a corresponding limited sampling of the spectral break. This source also has -1.2 which is redder than the UV slopes measured in spectroscopically confirmed objects at (Roberts-Borsani et al. 2024b, a), with the only notable exception of the X-ray emitting AGN GHZ9 at z=10.145 (=-1.10.12, Napolitano et al. 2024).
5 Constraints on the UV Luminosity Function beyond z=15
We explore in this section the implications that our findings may have on the evolution of the Luminosity Function (LF) at . In order to take into account the caveats discussed above on the potential contamination from rare classes of interlopers, we consider two opposite scenarios regarding the reliability of our F200W-dropout sample. In Case 1, we assume that all our 9 candidates are at , as indicated by the selection function of Fig. 1. As we have described above, all our candidates are selected following a standard, self-consistent approach which has proved to be effective at lower redshifts. Despite the concerns described above, we stress that all our candidates are at least consistent with the expected properties of galaxies in this redshift range, and albeit contamination is expected both on the basis of our simulations and selection results (Sect. 3), there are no compelling reasons to reject them a priori. In the opposite Case 2 we assume that all candidates are interlopers, corresponding to a non-detection of sources in every field, and compute the corresponding upper limits on the LF.
5.1 The computation of the LF
The LF has been computed following a standard approach described in detail in Castellano et al. (2023), which takes into account incompleteness and selection effects through imaging simulations that have been performed separately for each of the fields. Briefly, we inserted in blank regions of the observed images mock Lyman-break galaxies at and with a uniform distribution at mag. The observed magnitudes are obtained by randomly associating a model from a library based on BC03 models with metallicity , mag and a constant star-formation history. We assume that objects follow a circular Sersic (1968) light profile with index . Considering the lack of estimates of the size distribution at these redshifts we have assumed a fixed size of 0.2 kpc, which is consistent with the typical Rh of our candidates. In order to avoid overcrowding, simulations are performed by inserting 500 objects each time. Detection, photometry, and color selection on the simulated galaxies are performed in the same way as for the real catalogs. The simulated populations are used to estimate the completeness of our colour selections in each of the considered magnitude bins, hence the effective volume accessible in each field. In the case of A2744, which is affected by lensing, we adopt the approach in Eq. 1 of Castellano et al. (2023), namely the effective volumes in each bin are obtained by taking into account the area at different magnification levels computed on the basis of the model by Bergamini et al. (2023), and the relevant completeness for the selection of objects with the considered UV rest-frame magnitudes. While we remark that no candidates are found in this field, we limit our LF analysis to regions with , to avoid the small strongly lensed regions where systematic uncertainties may be significant, and source multiplicity would need to be taken into account in the simulation process.
We underline that our procedure for estimating the binned LF does not attempt to include the effect of contamination by leveraging the information contained in to weight the number of observed objects (see e.g. Donnan et al. 2024). However, for the reasons described above, we believe that the knowledge of the population of potential interlopers in LBG selections at these extreme redshifts is too uncertain at the moment, and we prefer to bracket the various options with the two opposite scenarios described above.
The results are reported in Tables 2 (z=15-20) and 3 (z=20-30), and shown in Fig. 8 and 9, respectively.
The comparison to available measurements of the LF at clearly shows that if all our 9 F200W-dropout candidates are at (Case 1), the LF continues a trend of very slow evolution, similar to recent findings at (Harikane et al. 2025; Donnan et al. 2024; Robertson et al. 2024; Casey et al. 2024; Finkelstein et al. 2024). Our estimates would be in agreement with the measurements at z=16-18 by Harikane et al. (2024), Whitler et al. (2025), Kokorev et al. (2024) and Pérez-González et al. (2025) in the same luminosity range. Most importantly, a similar scenario remains valid if we consider partial, but not complete, contamination of our sample and/or successful confirmation of some candidates in our extended sample. In fact, every point in the LF is originated by a small number of observed galaxies falling in that luminosity bin (typically 1 or 2), such that even if a small fraction of candidates is indeed at , the corresponding density in the considered bin would be comparable to the available estimates at 14. In addition, three of our candidates are brighter than the highest-redshift secure galaxy JADES-GS-z14-0 (MUV=-20.81, Carniani et al. 2024a): if confirmed, they would imply a number density at M-21 higher than current measurements at 14.
Needless to say, if we assume that none of our candidates are genuinely at (Case 2), the upper limits simply indicate that the actual LF is located at lower densities and luminosities, to an extent that we cannot establish with the existing data. The reader should not get confused by the fact that the upper limits of this scenario are not significantly different from the positive detections. The 1 upper limit in case of non-detection, computed on the basis of small number Poisson statistics (Gehrels 1986), corresponds to , a value very close to the observed densities in ?Case 1?.
The estimate of the galaxy number density at , for which no reference F277W-dropout candidates have been selected, yields a similar conclusion. The derived upper limits are consistent with the z25 estimates by Pérez-González et al. (2025), and somewhat above the existing measurements at and at , due to the reduced effective volume sampled by our surveys. This result demonstrates that significantly wider/deeper areas are in principle necessary to sample this redshift range - a point that will be further expanded below.
| (Case 1) | (Case 2) | ||
| Mpc-3 mag-1 | |||
| -22.5 | 1 | 0.06 | 0.12 |
| -21.5 | 2 | 0.14 | 0.13 |
| -20.5 | 4 | 0.50 | 0.23 |
| -19.5 | 2 | 1.2 | 1.1 |
| -18.5 | 0 | 5.4 | 5.4 |
a) The number of sources in the reference F200W-dropout sample in each rest-frame bin, and the resulting number densities assuming they are all at (Case 1) or that they are all interlopers (Case 2).
| Mpc-3 mag-1 | |
| -22.5 | 0.14 |
| -21.5 | 0.19 |
| -20.5 | 0.55 |
| -19.5 | 2.8 |
| -18.5 | 9.5 |
5.2 Comparison to theoretical predictions
We explore the implications of our results by showing in Fig. 10 and Fig. 11 a comparison with a variety of theoretical models aimed at explaining the mild evolution of the UV LF beyond z10.
5.2.1 Empirical models
To put our results in context, we first present a simple comparison with an empirically-adjusted, theoretically-motivated LF. The model is obtained starting from a standard calculation of the Cold Dark Matter Halo Mass Function (CDM-HMF). We adopt the Sheth & Tormen (1999) form, assuming a CDM linear power spectrum. Compared to other expressions proposed so far for the halo mass function (e.g., Yung et al. 2024b), this form provides the most extended high-mass tail and thus constitutes the most conservative form for our goals. This HMF is converted directly into a UV LF assuming the conversion curve at by Mason et al. (2015), brightened by exactly 1 magnitude to broadly match the observed UV LF by Donnan et al. (2024). We remark that we make no effort to physically motivate this brightening, which can be ascribed to a number of effects. We use it simply as a reference point to illustrate the evolution of the UV LF at higher redshifts under simple assumptions.
We let then evolve up to the LF at under the assumption of a non-evolving , which is shown in the same panel of Fig. 10. In practice, the entire evolution of the LF is driven by the corresponding evolution of the Press & Schechter HMF. As can be seen, the resulting evolution is extremely accelerated beyond , with a drop of up to two orders of magnitude at from to , and four orders of magnitudes up to , or, equivalently, by a drop in luminosity at constant density of 2 magnitudes from to and more than 4 mags from to . By construction, this reflects the evolution of the critical mass for collapse in the standard –CDM model, that indeed evolves dramatically at these redshifts (e.g., Menci et al. 2024). While the assumption of a constant beyond is certainly coarse and inadequate, it is a useful exercise to demonstrate how hard it is to imagine physical mechanisms that may effectively compensate for this fast evolution and maintain the UV LF significantly higher.
As an opposite case, we build an empirical model explicitly aimed at maximising the abundance of high-redshift galaxies under the extreme assumption that all baryons accreted onto a DM halo are instantaneously converted into stars (?max SFE? model). Specifically, we assume that the star formation rate equals the baryonic mass growth , where is the baryon mass fraction and is the dark-matter mass growth rate. The latter is computed after the fitting formula (based on N-body simulations) given in Correa et al. (2015), which depends on the halo mass and on the redshift. The UV luminosity associated with the different halo masses is then computed by assuming a star-formation efficiency =1 and a ratio of a dust-free template with metallicity Z=0.02 Z⊙, age=10 Myr, and a Chabrier (2003) IMF. Notably, the resulting UV LFs at 15 fall below our measurements (Fig. 10, right panel). To further include effects that maximize the UV luminosity associated with a given dark matter halo, we also allowed for a stochastic fluctuation of the star formation rate, adopting the simplified description proposed by Kravtsov & Belokurov (2024). In this approach the star formation rate is multiplied by , where is a correlated random number drawn from a Gaussian distribution with zero mean. Since our aim is not to provide a best-fit of the LFs but rather to derive a maximal UV luminosity associated with dark matter halos, we adopt a simplified treatment, where - instead of extracting from a proper distribution - we assume for it a fixed value . This is larger than the typical range for the rms value resulting from the analysis of stochasticity of the star formation rate in the high-resolution zoom-in simulations by Kravtsov & Belokurov (2024), and definitely larger than the value they assume in their best-fitting models for the UV luminosity function. Only the additional effect of extreme stochasticity, highlighted by a shaded region in Fig. 10, allows this simple empirical model to match the number density inferred from our reference F200W-dropout sample (consistently with the results by Pallottini & Ferrara 2023). A similar result may be obtained by allowing for top-heavy, or flat IMFs, which can increase by up to a factor of 10 compared to a Chabrier IMF.
These simple tests suggest that any smoothly evolving theoretical extrapolation of the UV LF would predict an evolution at stronger than our estimates, and that additional physical mechanisms must be at play to match such a high abundance of bright galaxies.
5.2.2 Analytic and semi-analytic models
A number of self-consistent physical models have been explored to understand the high-abundance of bright galaxies observed at z9 by JWST. In the ?Attenuation-free model? (AFM, Ferrara et al. 2023; Ziparo et al. 2023; Fiore et al. 2023; Ferrara et al. 2025a) radiation-driven outflows expel or lift the previously formed dust, thus boosting the UV luminosity to an extent that matches the observed LFs at z10-14. In the ?Feedback-free starbursts? scenario (FFB, Dekel et al. 2023; Li et al. 2024) the excess of bright galaxies is explained as the result of high densities and low metallicities yielding a extremly high star-formation efficiency at cosmic dawn. The DELPHI semi-analytic model (SAM) based on cold gas fractions and star formation efficiencies sampled from the sphinx simulations (Mauerhofer et al. 2025) explored two different mechanisms boosting the abundance of galaxies at 9: a stellar initial mass function (IMF) that becomes increasingly top-heavy with decreasing metallicity and increasing redshift (eIMF model), and star formation efficiencies that increase with increasing redshift (eSFE model). Similarly, the CAT SAM invokes a gradual transition in the IMF, modulated by metallicity and redshift to match the UV LFs at very high-redshift (Trinca et al. 2024). Finally, the recently updated Santa Cruz semi-analytic model (Somerville et al. in prep., see also Yung et al. 2024a) was run on dark matter halo merger trees extracted from the GUREFT simulations (Yung et al. 2024b), and incorporates a star formation efficiency that increases with increasing gas surface density, motivated by results from molecular cloud-scale simulations with radiative transfer. As overall galaxy surface densities are naturally predicted to be higher at early times in the CDM picture, these models predict higher star formation efficiencies and therefore more UV luminous galaxies at early times. The free parameter represents the fraction of the ISM that is in dense, star forming clouds.
Consistently with the simple empirical predictions described above, all these theoretical scenarios point towards a strong evolution at (Fig. 11). Both the AFM model, and the FFB one with star-formation efficiency , provide a good match to the z12 UV LF but predict an abundance of galaxies at z15 lower than our Case 1 LF. Similarly, the CAT SAM based on a standard IMF (Trinca et al. 2024), the DELPHI ?fiducial? model, and the Santa Cruz SAM with =0.1, fall below our estimates.
Interestingly, all models are inconsistent at 2 with the Case 1 abundance at M-20, but the tension is alleviated when assuming a change in physical properties at z15. The DELPHI eIMF and eSFE models, and the CAT model assuming an evolving IMF, are partially consistent with our estimates at -18.5–19.5. An evolution of the dense gas fraction from 0.1 at z=13 to 0.5 could match our estimates at all luminosities according to the Santa Cruz SAM. The FFB model with is also consistent with our Case 1 results at -20, but lower SFEs may also be viable if galaxies at 15 follow a relation between stellar mass and UV luminosity different than the z10 one adopted by Li et al. (2024).
These comparisons help us to put the results presented in this paper in context, leading to our main conclusions:
If even a fraction of the candidates presented here is indeed at , the tension with existing theoretical models, would be significant. In particular, the confirmation of bright (M-21) candidates would require deep revisions of our theoretical framework. A high abundance at the bright-end would imply a SFE close to 100%, or a substantial contribution from AGN or other very luminous sources, such as black holes (Pacucci et al. 2022) or primordial black holes (Matteri et al. 2025). In addition, a successful confirmation of any of the candidates in the extended F277W-dropout sample (Sect A) would imply a dramatic discrepancy with all theoretical models.
If instead all these candidates are interlopers, we are forced to conclude that future surveys will need to cover much wider areas to secure the selection of bright galaxies significantly beyond and to test predictions of current theoretical models. We will further investigate this point in the next Section.
6 Designing a survey to break the z=15 barrier
We built on the analysis of the ASTRODEEP-JWST fields to constrain the requirements for future observations designed to individuate robust samples of galaxies at .
We first discuss how to improve the robustness of candidate selection. We use the properties of potential, rare low-redshift contaminants discussed in Sect. 4.2 to constrain the relative depth between different JWST bands capable of discriminating between interlopers and genuine high-redshift galaxies. We note that in the surveys analysed here the bands immediately blueward of the Lyman break, i.e. F150W and F200W, are typically shallower by 0.5 mag and up to 1 mag than the LW bands F356W and F444W sampling the UV continuum up to z30. We have thus used the expected fluxes of all templates at that provide a good fit to our candidates to build mock JWST catalogs varying the relative depth between the various filters. We consider templates of the low-redshift solutions of all our samples, i.e. the reference F200W dropout objects, and the extended samples of F200W- and F277W-dropouts, and perturb the predicted fluxes with random Gaussian noise. We compared a reference scenario in which the NIRCam SW bands are 0.5 mags shallower than the LW ones, similarly to the observed fields, to scenarios in which the SW bands are as deep or deeper than the LW ones. For simplicity, we normalize the templates to have mag=29 in the F356W band and fix a depth 29AB at SNR=10 in the same band in all our simulations.
As expected, we find that the bands immediately redward of the expected break are crucial for discriminating between low- and high-redshift solutions. Increasing the depth of the F090W and/or F115W bands is basically ineffective in reducing the contamination rate, while having both F150W and the F200W as deep as the LW bands reduces by a factor of 3 the fraction of templates contaminating the colour selections. The best advantage is obtained when the two bands are 0.5 mag deeper than the LW ones at fixed SNR. In such a case, the contamination fraction is reduced by a factor of 10. Most importantly, such a deeper imaging at 1.5-2 would make it possible to adopt the more inclusive color threshold (z15-20, see Fig. 1) with a contamination rate of 0.5% of the considered templates, with a significant gain in the accessible color space.
We then addressed the question of the survey area required to detect a given number of sources as predicted by the various theoretical models discussed in the previous section. The results are shown in Fig. 12. Finding at least one object at the very bright end (M-20) of the UV LF at requires reaching a continuum magnitude around 28 (at SNR=10) over an area of sq.arcmin in most of the scenarios. Consistently with Figure 11, only models with extreme star-formation efficiency predict the detection of one object on areas comparable to our current data set ( sq.arcmin). Conversely, if one aims at the faint side of the LF (M-17) at , an area of sq. arcmin at a depth around is required to detect one object. Following our tests, in both cases observations at least mags deeper are required in the F150W and F200W bands to perform an unambiguous selection.
With these numbers in mind, it is intriguing to estimate the JWST observing time needed to reach these combinations of depth and size. For the bright side (M-20), according to the JWST exposure time calculator, NIRCam can observe at a depth of at SNR=10 both the F150W, F200W bands with a total of 2.5 hours of net exposure time per pointing. The simultaneous observation in two channels allows to observe F277W, F356W, and F444W at (SNR=10). An area of 12000 sq. arcmin which, according to most of the predictions, enables the detection of at least 2-3 sources with M-20, would require an investment of no less than 3000 hours plus overheads. Similarly, a deep pencil-beam NIRCam pointing in the F150W, F200W, F277W and F356W bands to unambiguously detect at SNR=10 at least two ultra-faint objects of continuum mag AB requires more than 1300 hours of net exposure. While this simple exercise is only meant to provide an order-of-magnitude estimate of the time needed for robust photometric selections, it clearly highlights that a thorough characterization of the available candidates and of the potential contaminants shall be considered a prerequisite for any future effort in this direction.
7 Summary and conclusions
The high abundance of galaxies beyond z10 can potentially be explained by several competing scenarios. Extending the constraints on the UV LF to the poorly explored range at z15 allows discriminating among different theoretical models, and individuating bright galaxies which are crucial to expand our knowledge on the first phases of galaxy formation. To this aim, we have analysed the ASTRODEEP-JWST photometric sample by Merlin et al. (2024), which provides consistent measurements on the major JWST deep surveys, to select bright galaxy candidates at z15-30. On the basis of mock observations mimicking the properties of our dataset, we have designed specific renditions of the Lyman-break selection technique that efficiently identify galaxy candidates in the redshift ranges and .
We isolated nine candidates at , while no objects are found at . A closer inspection of the selected candidates shows that despite exhibiting a 1.5 mag break, the selected objects consistently display multimodal redshift probability distributions across different SED-fitting codes and methodologies. The alternative solutions cover regions in the versus and versus planes different from the general populations at the same redshifts. Most importantly, they correspond to populations of low-mass () quiescent or dusty galaxies that have not been thoroughly investigated so far. This result is corroborated by the spectral properties of object UDS_56824, a confirmed interloper of our F200W-dropout selection, which is found to be a dusty () starburst galaxy at with mass . These results imply that while our candidates are, in principle, credible objects at , none of them would pass stringent selection criteria based on between different redshift solutions. In addition, considering that the low-redshift templates populate a basically unexplored parameter space whose galaxy density is unconstrained, using their as a weight to estimate the high-redshift UV LF appears to be an approach rife with uncertainties.
We adopted a more pragmatic approach of estimating the UV LF at z15-20 by assuming different contamination levels. The UV LF based on the, admittedly extreme, assumption of negligible contamination indicates a very mild evolution compared to estimates at immediately lower redshifts, at odds with all theoretical predictions. In particular, the confirmation of bright (M-21) candidates would require deep revisions of our theoretical framework, and might be even in contrast with any plausible model under a standard –CDM cosmological scenario. However, the tension with theoretical models is so significant that even if only a small fraction of the candidates is confirmed to be at , a further evolution of physical properties, such as IMF or star-formation efficiency would be required to explain the observed number densities. If instead all the analysed candidates are interlopers, we are forced to conclude that future surveys will need to cover much wider areas to secure the selection of bright galaxies significantly beyond .
According to a variety of theoretical models, finding at least one object at the very bright end (M-20) of the UV LF requires surveying very large areas, ranging from 500 sq. arcmin. to 1000 sq. arcmin at , and to more than 2000 sq. arcmin at 20. However, our analysis shows that a large area coverage is not the only required ingredient, because the depth achieved by current surveys is not optimal to avoid contamination. A simple test based on the properties of the low-redshift solutions of our candidates indicates that NIRCam imaging in the F150W and F200W bands should be at least as deep as the observations in NIRCam LW bands, and possibly 0.5 mag deeper, to decrease the contamination rate significantly. This is a demanding requirement for both ultra-deep pencil beam observations targeting the LF faint-end and for large area surveys sampling the bright-end. We argue that a more pragmatic approach should aim, first of all, at a thorough spectroscopic characterization of the candidates available on current surveys and of potential contaminating populations to gather key information to plan future surveys aimed at breaking the current redshift records.
Acknowledgements.
We thank J. Dunlop and R. Ellis for the useful discussions. We acknowledge financial support from NASA through grant JWST-ERS-1324. Support was also provided by the PRIN 2022 MUR project 2022CB3PJ3 – First Light And Galaxy aSsembly (FLAGS) funded by the European Union – Next Generation EU, and by INAF Mini-grant “Reionization and Fundamental Cosmology with High-Redshift Galaxies.” P.G.P.-G. and L.C. acknowledge support from grant PID2022-139567NB-I00 funded by Spanish Ministerio de Ciencia e Innovación MCIN/AEI/10.13039/501100011033, FEDER Una manera de hacer Europa. L.N. acknowledges support from grant “Progetti per Avvio alla Ricerca - Tipo 1, Unveiling Cosmic Dawn: Galaxy Evolution with CAPERS” (AR1241906F947685). RA acknowledges financial support from grants PID2023-147386NB-I00 ”XTREM” and PID2022-136598NB-C32 “Estallidos8”, funded by MCIN/AEI/10.13039/501100011033 and ERDF/EU, and the State Agency for Research of the Spanish MCIU through ‘Center of Excellence Severo Ochoa’ award to the IAA-CSIC (SEV-2017-0709) and CEX2021-001131-S. F.P. acknowledges support from a Clay Fellowship administered by the Smithsonian Astrophysical Observatory. The Flatiron Institute is supported by the Simons Foundation. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs #1324, 1345, 1180, 1210, 1837, 2561, 2756, 3990, and 6368.References
- Algera et al. (2024) Algera, H. S. B., Inami, H., De Looze, I., et al. 2024, MNRAS, 533, 3098
- Arrabal Haro et al. (2023a) Arrabal Haro, P., Dickinson, M., Finkelstein, S. L., et al. 2023a, ApJ, 951, L22
- Arrabal Haro et al. (2023b) Arrabal Haro, P., Dickinson, M., Finkelstein, S. L., et al. 2023b, Nature, 622, 707
- Austin et al. (2023) Austin, D., Adams, N., Conselice, C. J., et al. 2023, ApJ, 952, L7
- Bagley et al. (2024) Bagley, M. B., Pirzkal, N., Finkelstein, S. L., et al. 2024, ApJ, 965, L6
- Bergamini et al. (2023) Bergamini, P., Acebron, A., Grillo, C., et al. 2023, arXiv e-prints, arXiv:2303.10210
- Bertin & Arnouts (1996) Bertin, E. & Arnouts, S. 1996, A&AS, 117, 393
- Bezanson et al. (2022) Bezanson, R., Labbe, I., Whitaker, K. E., et al. 2022, arXiv e-prints, arXiv:2212.04026
- Birrer et al. (2021) Birrer, S., Shajib, A., Gilman, D., et al. 2021, The Journal of Open Source Software, 6, 3283
- Bisigello et al. (2025) Bisigello, L., Gandolfi, G., Feltre, A., et al. 2025, A&A, 693, L18
- Bisigello et al. (2023) Bisigello, L., Gandolfi, G., Grazian, A., et al. 2023, A&A, 676, A76
- Boquien et al. (2019) Boquien, M., Burgarella, D., Roehlly, Y., et al. 2019, A&A, 622, A103
- Bouwens et al. (2020) Bouwens, R., González-López, J., Aravena, M., et al. 2020, ApJ, 902, 112
- Bouwens et al. (2021) Bouwens, R. J., Oesch, P. A., Stefanon, M., et al. 2021, AJ, 162, 47
- Brammer et al. (2008) Brammer, G. B., van Dokkum, P. G., & Coppi, P. 2008, ApJ, 686, 1503
- Bruzual & Charlot (2003) Bruzual, G. & Charlot, S. 2003, MNRAS, 344, 1000
- Calzetti et al. (2000) Calzetti, D., Armus, L., Bohlin, R. C., et al. 2000, ApJ, 533, 682
- Carnall et al. (2024) Carnall, A. C., Cullen, F., McLure, R. J., et al. 2024, MNRAS, 534, 325
- Carnall et al. (2019) Carnall, A. C., McLure, R. J., Dunlop, J. S., et al. 2019, MNRAS, 490, 417
- Carnall et al. (2018) Carnall, A. C., McLure, R. J., Dunlop, J. S., & Davé, R. 2018, MNRAS, 480, 4379
- Carniani et al. (2024a) Carniani, S., D’Eugenio, F., Ji, X., et al. 2024a, arXiv e-prints, arXiv:2409.20533
- Carniani et al. (2024b) Carniani, S., Hainline, K., D’Eugenio, F., et al. 2024b, Nature, 633, 318
- Casey et al. (2024) Casey, C. M., Akins, H. B., Shuntov, M., et al. 2024, ApJ, 965, 98
- Casey et al. (2019) Casey, C. M., Zavala, J. A., Aravena, M., et al. 2019, ApJ, 887, 55
- Castellano et al. (2023) Castellano, M., Fontana, A., Treu, T., et al. 2023, ApJ, 948, L14
- Castellano et al. (2022) Castellano, M., Fontana, A., Treu, T., et al. 2022, ApJ, 938, L15
- Castellano et al. (2024) Castellano, M., Napolitano, L., Fontana, A., et al. 2024, ApJ, 972, 143
- Castellano et al. (2014) Castellano, M., Sommariva, V., Fontana, A., et al. 2014, A&A, 566, A19
- Chabrier (2003) Chabrier, G. 2003, PASP, 115, 763
- Cole et al. (2025) Cole, J. W., Papovich, C., Finkelstein, S. L., et al. 2025, ApJ, 979, 193
- Conselice et al. (2024) Conselice, C. J., Adams, N., Harvey, T., et al. 2024, arXiv e-prints, arXiv:2407.14973
- Correa et al. (2015) Correa, C. A., Wyithe, J. S. B., Schaye, J., & Duffy, A. R. 2015, MNRAS, 450, 1514
- Dale et al. (2014) Dale, D. A., Helou, G., Magdis, G. E., et al. 2014, ApJ, 784, 83
- de Graaff et al. (2025) de Graaff, A., Setton, D. J., Brammer, G., et al. 2025, Nature Astronomy, 9, 280
- Dekel et al. (2023) Dekel, A., Sarkar, K. C., Birnboim, Y., Mandelker, N., & Li, Z. 2023, MNRAS, 523, 3201
- Ding et al. (2020) Ding, X., Silverman, J., Treu, T., et al. 2020, ApJ, 888, 37
- Donnan et al. (2023) Donnan, C. T., McLeod, D. J., McLure, R. J., et al. 2023, MNRAS, 520, 4554
- Donnan et al. (2024) Donnan, C. T., McLure, R. J., Dunlop, J. S., et al. 2024, MNRAS, 533, 3222
- Dudzevičiūtė et al. (2020) Dudzevičiūtė, U., Smail, I., Swinbank, A. M., et al. 2020, MNRAS, 494, 3828
- Eisenstein et al. (2023) Eisenstein, D. J., Willott, C., Alberts, S., et al. 2023, arXiv e-prints, arXiv:2306.02465
- Ferland et al. (2013) Ferland, G. J., Porter, R. L., van Hoof, P. A. M., et al. 2013, Revista Mexicana de Astronomía y Astrofísica, 49, 137
- Ferrara et al. (2025a) Ferrara, A., Carniani, S., di Mascia, F., et al. 2025a, A&A, 694, A215
- Ferrara et al. (2023) Ferrara, A., Pallottini, A., & Dayal, P. 2023, MNRAS, 522, 3986
- Ferrara et al. (2025b) Ferrara, A., Pallottini, A., & Sommovigo, L. 2025b, A&A, 694, A286
- Finkelstein et al. (2025) Finkelstein, S. L., Bagley, M. B., Arrabal Haro, P., et al. 2025, arXiv e-prints, arXiv:2501.04085
- Finkelstein et al. (2022) Finkelstein, S. L., Bagley, M. B., Arrabal Haro, P., et al. 2022, ApJ, 940, L55
- Finkelstein et al. (2023) Finkelstein, S. L., Leung, G. C. K., Bagley, M. B., et al. 2023, arXiv e-prints, arXiv:2311.04279
- Finkelstein et al. (2024) Finkelstein, S. L., Leung, G. C. K., Bagley, M. B., et al. 2024, ApJ, 969, L2
- Fiore et al. (2023) Fiore, F., Ferrara, A., Bischetti, M., Feruglio, C., & Travascio, A. 2023, ApJ, 943, L27
- Fontana et al. (2000) Fontana, A., D’Odorico, S., Poli, F., et al. 2000, AJ, 120, 2206
- Gandolfi et al. (2022) Gandolfi, G., Lapi, A., Ronconi, T., & Danese, L. 2022, Universe, 8, 589
- Gandolfi et al. (2025) Gandolfi, G., Rodighiero, G., Bisigello, L., et al. 2025, arXiv e-prints, arXiv:2502.02637
- Geach et al. (2017) Geach, J. E., Dunlop, J. S., Halpern, M., et al. 2017, MNRAS, 465, 1789
- Gehrels (1986) Gehrels, N. 1986, ApJ, 303, 336
- Giavalisco (2002) Giavalisco, M. 2002, ARA&A, 40, 579
- Glazebrook et al. (2023) Glazebrook, K., Nanayakkara, T., Jacobs, C., et al. 2023, ApJ, 947, L25
- Glazebrook et al. (2024) Glazebrook, K., Nanayakkara, T., Schreiber, C., et al. 2024, Nature, 628, 277
- Grogin et al. (2011) Grogin, N. A., Kocevski, D. D., Faber, S. M., et al. 2011, ApJS, 197, 35
- Hainline et al. (2024) Hainline, K. N., Johnson, B. D., Robertson, B., et al. 2024, ApJ, 964, 71
- Harikane et al. (2025) Harikane, Y., Inoue, A. K., Ellis, R. S., et al. 2025, ApJ, 980, 138
- Harikane et al. (2024) Harikane, Y., Nakajima, K., Ouchi, M., et al. 2024, ApJ, 960, 56
- Harikane et al. (2023) Harikane, Y., Ouchi, M., Oguri, M., et al. 2023, ApJS, 265, 5
- Holwerda et al. (2018) Holwerda, B. W., Bridge, J. S., Ryan, R., et al. 2018, A&A, 620, A132
- Holwerda et al. (2024) Holwerda, B. W., Hsu, C.-C., Hathi, N., et al. 2024, MNRAS, 529, 1067
- Hutter et al. (2025) Hutter, A., Cueto, E. R., Dayal, P., et al. 2025, A&A, 694, A254
- Juneau et al. (2014) Juneau, S., Bournaud, F., Charlot, S., et al. 2014, ApJ, 788, 88
- Killi et al. (2024) Killi, M., Watson, D., Brammer, G., et al. 2024, A&A, 691, A52
- Koekemoer et al. (2011) Koekemoer, A. M., Faber, S. M., Ferguson, H. C., et al. 2011, ApJS, 197, 36
- Kokorev et al. (2024) Kokorev, V., Atek, H., Chisholm, J., et al. 2024, arXiv e-prints, arXiv:2411.13640
- Kravtsov & Belokurov (2024) Kravtsov, A. & Belokurov, V. 2024, arXiv e-prints, arXiv:2405.04578
- Kron (1980) Kron, R. G. 1980, ApJS, 43, 305
- Larson et al. (2022) Larson, R. L., Hutchison, T. A., Bagley, M., et al. 2022, arXiv e-prints, arXiv:2211.10035
- Leung et al. (2023) Leung, G. C. K., Bagley, M. B., Finkelstein, S. L., et al. 2023, ApJ, 954, L46
- Li et al. (2024) Li, Z., Dekel, A., Sarkar, K. C., et al. 2024, A&A, 690, A108
- Marley et al. (2021) Marley, M. S., Saumon, D., Visscher, C., et al. 2021, ApJ, 920, 85
- Mason et al. (2015) Mason, C. A., Trenti, M., & Treu, T. 2015, ApJ, 813, 21
- Matteri et al. (2025) Matteri, A., Ferrara, A., & Pallottini, A. 2025, arXiv e-prints, arXiv:2503.18850
- Mauerhofer et al. (2025) Mauerhofer, V., Dayal, P., Haehnelt, M. G., et al. 2025, arXiv e-prints, arXiv:2502.02647
- McGaugh et al. (2024) McGaugh, S. S., Schombert, J. M., Lelli, F., & Franck, J. 2024, ApJ, 976, 13
- McLeod et al. (2024) McLeod, D. J., Donnan, C. T., McLure, R. J., et al. 2024, MNRAS, 527, 5004
- McLure et al. (2018) McLure, R. J., Dunlop, J. S., Cullen, F., et al. 2018, MNRAS, 476, 3991
- Menci et al. (2024) Menci, N., Sen, A. A., & Castellano, M. 2024, ApJ, 976, 227
- Merlin et al. (2022) Merlin, E., Bonchi, A., Paris, D., et al. 2022, arXiv e-prints, arXiv:2207.11701
- Merlin et al. (2019) Merlin, E., Pilo, S., Fontana, A., et al. 2019, A&A, 622, A169
- Merlin et al. (2024) Merlin, E., Santini, P., Paris, D., et al. 2024, A&A, 691, A240
- Mitsuhashi et al. (2024) Mitsuhashi, I., Harikane, Y., Bauer, F. E., et al. 2024, ApJ, 971, 161
- Morishita et al. (2024) Morishita, T., Mason, C. A., Kreilgaard, K. C., et al. 2024, arXiv e-prints, arXiv:2412.04211
- Naidu et al. (2022a) Naidu, R. P., Matthee, J., Oesch, P. A., et al. 2022a, MNRAS, 510, 4582
- Naidu et al. (2022b) Naidu, R. P., Oesch, P. A., Setton, D. J., et al. 2022b, arXiv e-prints, arXiv:2208.02794
- Napolitano et al. (2025) Napolitano, L., Castellano, M., Pentericci, L., et al. 2025, A&A, 693, A50
- Napolitano et al. (2024) Napolitano, L., Castellano, M., Pentericci, L., et al. 2024, arXiv e-prints, arXiv:2410.18763
- Oesch et al. (2023) Oesch, P. A., Brammer, G., Naidu, R. P., et al. 2023, MNRAS, 525, 2864
- Oke & Gunn (1983) Oke, J. B. & Gunn, J. E. 1983, ApJ, 266, 713
- Ono et al. (2023) Ono, Y., Harikane, Y., Ouchi, M., et al. 2023, ApJ, 951, 72
- Ono et al. (2025) Ono, Y., Ouchi, M., Harikane, Y., et al. 2025, arXiv e-prints, arXiv:2502.08885
- Pacucci et al. (2022) Pacucci, F., Dayal, P., Harikane, Y., Inoue, A. K., & Loeb, A. 2022, MNRAS, 514, L6
- Pacucci et al. (2023) Pacucci, F., Nguyen, B., Carniani, S., Maiolino, R., & Fan, X. 2023, ApJ, 957, L3
- Pallottini & Ferrara (2023) Pallottini, A. & Ferrara, A. 2023, A&A, 677, L4
- Paris et al. (2023) Paris, D., Merlin, E., Fontana, A., et al. 2023, ApJ, 952, 20
- Peng et al. (2010) Peng, C. Y., Ho, L. C., Impey, C. D., & Rix, H.-W. 2010, AJ, 139, 2097
- Pérez-González et al. (2023a) Pérez-González, P. G., Barro, G., Annunziatella, M., et al. 2023a, ApJ, 946, L16
- Pérez-González et al. (2024a) Pérez-González, P. G., Barro, G., Rieke, G. H., et al. 2024a, ApJ, 968, 4
- Pérez-González et al. (2023b) Pérez-González, P. G., Costantin, L., Langeroodi, D., et al. 2023b, arXiv e-prints, arXiv:2302.02429
- Pérez-González et al. (2024b) Pérez-González, P. G., D‘Eugenio, F., Rodríguez del Pino, B., et al. 2024b, arXiv e-prints, arXiv:2405.03744
- Pérez-González et al. (2025) Pérez-González, P. G., Östlin, G., Costantin, L., et al. 2025, arXiv e-prints, arXiv:2503.15594
- Pérez-González et al. (2024c) Pérez-González, P. G., Rinaldi, P., Caputi, K. I., et al. 2024c, ApJ, 969, L10
- Rinaldi et al. (2024) Rinaldi, P., Navarro-Carrera, R., Caputi, K. I., et al. 2024, arXiv e-prints, arXiv:2406.13554
- Roberts-Borsani et al. (2024a) Roberts-Borsani, G., Bagley, M., Rojas-Ruiz, S., et al. 2024a, arXiv e-prints, arXiv:2407.17551
- Roberts-Borsani et al. (2024b) Roberts-Borsani, G., Treu, T., Shapley, A., et al. 2024b, arXiv e-prints, arXiv:2403.07103
- Robertson et al. (2024) Robertson, B., Johnson, B. D., Tacchella, S., et al. 2024, ApJ, 970, 31
- Rodighiero et al. (2023) Rodighiero, G., Bisigello, L., Iani, E., et al. 2023, MNRAS, 518, L19
- Santini et al. (2021) Santini, P., Castellano, M., Merlin, E., et al. 2021, A&A, 652, A30
- Santini et al. (2023) Santini, P., Fontana, A., Castellano, M., et al. 2023, ApJ, 942, L27
- Schaerer & de Barros (2009) Schaerer, D. & de Barros, S. 2009, A&A, 502, 423
- Schreiber et al. (2017) Schreiber, C., Elbaz, D., Pannella, M., et al. 2017, A&A, 602, A96
- Sersic (1968) Sersic, J. L. 1968, Atlas de galaxias australes (Cordoba, Argentina: Observatorio Astronomico, 1968)
- Sheth & Tormen (1999) Sheth, R. K. & Tormen, G. 1999, MNRAS, 308, 119
- Stanway & Eldridge (2018) Stanway, E. R. & Eldridge, J. J. 2018, MNRAS, 479, 75
- Tacchella et al. (2023) Tacchella, S., Eisenstein, D. J., Hainline, K., et al. 2023, ApJ, 952, 74
- Treu et al. (2022) Treu, T., Roberts-Borsani, G., Bradac, M., et al. 2022, ApJ, 935, 110
- Trinca et al. (2024) Trinca, A., Schneider, R., Valiante, R., et al. 2024, MNRAS[arXiv:2305.04944]
- Vulcani et al. (2017) Vulcani, B., Trenti, M., Calvi, V., et al. 2017, ApJ, 836, 239
- Weibel et al. (2024) Weibel, A., de Graaff, A., Setton, D. J., et al. 2024, arXiv e-prints, arXiv:2409.03829
- Westcott et al. (2024) Westcott, L., Conselice, C. J., Harvey, T., et al. 2024, arXiv e-prints, arXiv:2412.14970
- Whitler et al. (2025) Whitler, L., Stark, D. P., Topping, M. W., et al. 2025, arXiv e-prints, arXiv:2501.00984
- Williams et al. (2018) Williams, C. C., Curtis-Lake, E., Hainline, K. N., et al. 2018, ApJS, 236, 33
- Yan et al. (2023a) Yan, H., Ma, Z., Ling, C., Cheng, C., & Huang, J.-S. 2023a, ApJ, 942, L9
- Yan et al. (2023b) Yan, H., Sun, B., Ma, Z., & Ling, C. 2023b, arXiv e-prints, arXiv:2311.15121
- Yung et al. (2024a) Yung, L. Y. A., Somerville, R. S., Finkelstein, S. L., Wilkins, S. M., & Gardner, J. P. 2024a, MNRAS, 527, 5929
- Yung et al. (2024b) Yung, L. Y. A., Somerville, R. S., Nguyen, T., et al. 2024b, MNRAS, 530, 4868
- Zavala et al. (2023) Zavala, J. A., Buat, V., Casey, C. M., et al. 2023, ApJ, 943, L9
- Ziparo et al. (2023) Ziparo, F., Ferrara, A., Sommovigo, L., & Kohandel, M. 2023, MNRAS, 520, 2445
Appendix A The extended samples of z15 dropout candidates
We present here the extended samples of z (ID, coordinates and F356W magnitudes from M24 in Table 4, SED, thumbnails and in Fig. 13 and 14) and z (Table 5, Fig. 15 and 16) candidates. We remark that while these objects also have colours and consistent with high-redshift solutions, and shall be considered of interest for potential spectroscopic follow-up, the contamination level in these samples is expected to be significant (Sect. 3.1).
| ID | R.A. | Dec | F356W |
| deg. | deg. | AB | |
| A2744_36169 | 3.506835 | -30.303378 | 27.81 0.12 |
| COSMOS_378 | 150.139836 | 2.162604 | 27.59 0.15 |
| COSMOS_10730 | 150.111759 | 2.208056 | 27.66 0.22 |
| COSMOS_28354 | 150.076753 | 2.254899 | 27.14 0.13 |
| COSMOS_34866 | 150.153167 | 2.269104 | 28.66 0.14 |
| COSMOS_45778 | 150.109207 | 2.292050 | 28.45 0.17 |
| COSMOS_47748 | 150.155793 | 2.295937 | 27.93 0.21 |
| COSMOS_51984 | 150.124225 | 2.304599 | 28.90 0.16 |
| COSMOS_66011 | 150.168328 | 2.332822 | 29.16 0.20 |
| COSMOS_78212 | 150.155200 | 2.356336 | 27.83 0.15 |
| COSMOS_82058 | 150.088237 | 2.364684 | 28.77 0.23 |
| COSMOS_90343 | 150.166521 | 2.382355 | 27.66 0.13 |
| COSMOS_96354 | 150.149768 | 2.397964 | 29.04 0.22 |
| COSMOS_101131 | 150.143521 | 2.410134 | 27.89 0.14 |
| COSMOS_102115 | 150.192712 | 2.412584 | 28.90 0.26 |
| COSMOS_116352 | 150.170290 | 2.456889 | 28.54 0.24 |
| COSMOS_117020 | 150.145445 | 2.459396 | 28.56 0.19 |
| COSMOS_119125 | 150.184999 | 2.466924 | 28.04 0.22 |
| UDS_78048 | 34.397548 | -5.178046 | 26.83 0.10 |
| UDS_132278 | 34.354643 | -5.110434 | 27.15 0.11 |
| NGDEEP_3939 | 53.270238 | -27.861668 | 30.15 0.18 |
| JADES-GS_11943 | 53.030331 | -27.877975 | 29.42 0.11 |
| JADES-GN_1801 | 189.238294 | 62.148230 | 28.55 0.10 |
| JADES-GN_6483 | 189.324320 | 62.165237 | 28.02 0.12 |
| JADES-GN_14592 | 189.224961 | 62.188365 | 27.79 0.09 |
| JADES-GN_52334 | 188.989191 | 62.290761 | 27.54 0.10 |
| ID | R.A. | Dec | F356W |
| deg. | deg. | AB | |
| A2744_4252 | 3.658016 | -30.426589 | 27.49 0.12 |
| A2744_26717 | 3.500812 | -30.354774 | 29.89 0.20 |
| COSMOS_2139 | 150.132819 | 2.175271 | 27.97 0.19 |
| COSMOS_21874 | 150.077042 | 2.238654 | 28.56 0.19 |
| COSMOS_30664 | 150.109780 | 2.259779 | 26.70 0.16 |
| COSMOS_36047 | 150.178087 | 2.271629 | 28.38 0.19 |
| COSMOS_47136 | 150.133938 | 2.294643 | 29.32 0.29 |
| COSMOS_61893 | 150.065480 | 2.324540 | 26.36 0.18 |
| COSMOS_66588 | 150.122411 | 2.333940 | 29.75 0.27 |
| COSMOS_87695 | 150.129665 | 2.376409 | 28.62 0.10 |
| COSMOS_91135 | 150.183896 | 2.384041 | 28.57 0.16 |
| COSMOS_110332 | 150.191351 | 2.435514 | 27.96 0.13 |
| UDS_9701 | 34.261719 | -5.302606 | 26.85 0.08 |
| UDS_40510 | 34.388952 | -5.246670 | 27.49 0.09 |
| UDS_51790 | 34.247686 | -5.225198 | 27.25 0.08 |
| UDS_74000 | 34.268062 | -5.184911 | 27.48 0.08 |
| UDS_95077 | 34.345233 | -5.150057 | 28.24 0.11 |
| UDS_103005 | 34.402545 | -5.136976 | 27.63 0.14 |
| UDS_130776 | 34.434420 | -5.108762 | 28.35 0.14 |

### v2

@@
 Results.
-We identify nine candidates at , while no objects are found based on the color selection criteria.
+We identify five candidates at , while no objects are found based on the color selection criteria.
 Despite exhibiting a 1.5 mag break, all the objects display multimodal redshift probability distributions across different SED-fitting codes and methodologies.
 The alternative solutions correspond to poorly understood populations of low-mass quiescent or dusty galaxies at z3-7.
-This conclusion is supported by the analysis of a NIRSpec spectrum recently acquired by the CAPERS program for one interloper object, which is confirmed to be a dusty ( mag) starburst galaxy at .
+This conclusion is supported by the analysis of five F200W-dropout objects that we find to be interlopers on the basis of NIRSpec PRISM spectra: four dusty star-forming galaxies at z2.2-6.6, and a passive galaxy at z=4.91 with log 9.
 We measured the UV luminosity function under different assumptions on the contamination level within our sample.
@@
 2023a, b; Roberts-Borsani et al.
-2024b; Castellano et al.
+2024, 2025b; Castellano et al.
 2024; Napolitano et al.
 2024) and up to (Carniani et al.
-2024b, a), the current record holder.
+2024, 2025; Naidu et al.
+2025).
 This situation is in stark contrast with the poor constraints available at .
@@
 2023; Conselice et al.
-2024; Robertson et al.
+2025; Robertson et al.
 2024; Kokorev et al.
-2024; Whitler et al.
+2025; Whitler et al.
 2025; Gandolfi et al.
 2025; Pérez-González et al.
-2025), none of which has yet been confirmed spectroscopically.
+2025b), none of which has yet been confirmed spectroscopically.
 The most striking example is object CEERS-93316, which appeared to be a strong candidate based on CEERS imaging data (Donnan et al.
@@
 2023; Pérez-González et al.
-2023a, 2024c; Glazebrook et al.
+2023a, 2024b; Glazebrook et al.
 2023; Bisigello et al.
@@
 In fact, the different explanations that have been invoked to explain the mild evolution of the UV LF at z10 differ in their predictions at earliest times (Kokorev et al.
-2024; Pérez-González et al.
-2025).
+2025; Pérez-González et al.
+2025b).
 For instance, sustained high luminosity density beyond is favoured by changes in the initial mass function (IMF) or by an increased star-formation efficiency (e.g.
@@
 2025; Mauerhofer et al.
-2025), as well as by a rapid assembly of baryons (McGaugh et al.
+2025), as well as by a rapid assembly of baryons (Haslbauer et al.
+2022; McGaugh et al.
 2024), while the prediction of an earlier phase of dusty-enshrouded star-formation (e.g.
@@
 3 we describe our specific renditions of the Lyman-break technique and the results of our search for galaxy samples at and .
-We investigate in detail alternative low-redshift solutions of our candidates, and the spectroscopic properties of one newly confirmed interloper at z=6.56 in Sect.
+We investigate in detail alternative low-redshift solutions of our candidates, and the spectroscopic properties of five confirmed interlopers at z2-7 in Sect.
 4.
@@
 Labbé, Bezanson et al.
-2022), DDT 2756 (P.I.
+2024), DDT 2756 (P.I.
 Chen), and GO 3990 (P.I.
 Morishita, Morishita et al.
-2024).
+2025).
 The considered observations cover a significant range in both area and depth, from relatively wide surveys as PRIMER-UDS (250 sq.
@@
 2023).
+In the present paper we use total fluxes based on the optimal apertures as defined by M24, the signal-to-noise ratios are evaluated in 2 times the PSF full-width at half maximum (FWHM) for each band.
 The seven fields comprise a total of 531173 objects in an area of 615 sq.
@@
 For the present work we built the weighted average stacks of all the ACS and WFC3 HST bands available in each field, respectively.
-The SNR of all ASTRODEEP-JWST sources was measured within an aperture with a diameter of 2 times the PSF full-width at half maximum (FWHM) on both the HST stacks, and used as described in the following section to constrain non-detection blueward of the Lyman-break.
+The SNR of all ASTRODEEP-JWST sources was measured within an aperture with a diameter of 2 times the PSF FWHM on both the HST stacks, and used as described in the following section to constrain non-detection blueward of the Lyman-break.
 3 Lyman-break selection at z15-30
@@
 After excluding these two interlopers, we are left with an initial sample of 10 F200W dropouts.
-Of these, the only candidate selected in the PRIMER-UDS field (ID=56824 in M24) has been recently observed with NIRSpec PRISM observations by the CAPERS survey (GO-6368, P.I.
+Of these, five candidates (1 in UDS, 1 in CEERS and 3 in COSMOS) have been recently observed with NIRSpec PRISM by the CAPERS survey (GO-6368, P.I.
 M.
-Dickinson), yielding a spectroscopic redshift of , as discussed in more detail in Sect.
+Dickinson), finding that they are interlopers with a spectroscopic redshift in the range , as discussed in more detail in Sect.
 4.3.
 The final ?reference?
-sample of F200W dropout candidates that we shall discuss in this paper is therefore made of 9 sources.
-We performed additional checks on their reliability.
-First of all, we measured the SNR within an aperture of 2 times the PSF FWHM on a stack of the NIRCam F090W, F115W and F150W bands available for each of them, finding that they are all non-detected at SNR2.
-We then computed the SNR on the NIRCam F090W, F115W and F150W bands in two other apertures measured by M24, namely an aperture with a radius of 0.1 arcsec (R01), and the one with a diameter of 3 times the PSF-FWHM.
-All objects are non-detected in these bands at SNR2 in all cases, except COSMOS_35731 having SNR=2.1 in the F150W band in the R01 aperture (compared to SNR=1.4 and SNR1 in the 2FWHM and 3FWHM apertures, respectively).
-We estimated the MUV of the nine candidates by converting to the rest-frame the observed F277W magnitude assuming a redshift z=18 where our redshift selection function peaks, and their UV slope by fitting the F277W, F356W and F444W bands.
-We find three candidates brighter than M-21, with the brightest object in the sample COSMOS_107923 having M-22.7.
-Four of the objects have a red -2, the remaining ones being consistent with a flat or moderately blue UV slope ().
-We have measured their half-light radius by fitting the light distribution in the F277W band with GALIGHT (Ding et al.
-2020; Birrer et al.
-2021) assuming a Sersic (1968) profile with free index and fixing the redshift at the best-fit solution at .
-We find half-light radii consistent with those of observed in galaxies at z10-15 (Westcott et al.
-2024; Ono et al.
-2025).
-Our candidates have a typical half-light radius of 0.2-0.3 kpc, except COSMOS_31168, which is more extended (R0.5 kpc), and the compact source COSMOS_35731 (R1.5 pixel, amounting to 0.13 kpc) which is marginally resolved.
-Although its estimated size is comparable to that of other known sources at slightly lower redshifts (e.g., Ono et al.
-2023; Tacchella et al.
-2023) we verified both with GALIGHT and GALFIT (Peng et al.
-2010) that a fit with a PSF-like profile provides comparable and residuals as the Sersic one.
-We find a high field-to-field variance with 6 candidates found in PRIMER-COSMOS, 2 in CEERS, 1 in JADES-GN, and no sources selected in both JADES-GS and NGDEEP.
+sample of F200W dropout candidates that we shall discuss in this paper is therefore made of 5 sources: 3 candidates in PRIMER-COSMOS, 1 in CEERS, 1 in JADES-GN, and no sources selected in both JADES-GS and NGDEEP.
 As described above, neither A2744 nor PRIMER-UDS, which is the widest/shallowest among the considered fields, contribute to the final sample as the two selected sources are confirmed interlopers.
-Interestingly, candidates in both COSMOS and CEERS are close on the sky plane, such that if their high-redshift nature will be confirmed they would be at 1-5 physical Mpc distance from each other, possibly implying that they are part of distant overdensities.
-We do not find any object meeting our reference F277W-dropout criteria.
-In addition, we select extended samples of 26 F200W dropout objects, and of 19 F277W-dropout sources.
+Interestingly, candidates in COSMOS are close on the sky plane, such that if their high-redshift nature will be confirmed they would be at 1-5 physical Mpc distance from each other, possibly implying that they are part of distant overdensities.
+In addition, we select 26 F200W-dropout objects meeting the less conservative color thresholds described in Sec.
+3 (?extended sample?).
+We do not find any object meeting our reference F277W-dropout criteria, while 19 sources are included in the relevant ?extended sample?.
 We show in Fig.
@@
 The SEDs and main properties of the objects in the extended samples are presented in the Appendix A.
+We performed additional checks on the reliability of the five objects in the reference sample of F200W-dropout sources.
+First of all, we measured the SNR within an aperture of 2 times the PSF FWHM on a stack of the NIRCam F090W, F115W and F150W bands available for each of them, finding that they are all non-detected at SNR2.
+We then computed the SNR on the NIRCam F090W, F115W and F150W bands in two other apertures measured by M24, namely an aperture with a radius of 0.1 arcsec (R01), and the one with a diameter of 3 times the PSF-FWHM.
+All objects are non-detected in these bands at SNR2 in all cases.
+We then searched for MIRI imaging observations at the position of our candidates.
+The three COSMOS sources are covered by observations by the PRIMER and COSMOS-3D (GO-5893, PI K.
+Kakiichi) programs, while CEERS_17384 has been observed by the MIRI EGS Galaxy and AGN survey (MEGA; GO-3794, PI A.
+Kirkpatrick).
+We reduced the MIRI imaging datasets in all fields using the Rainbow JWST pipeline as described in Pérez-González et al.
+(2024a), and measured the SNR at the position of the candidates on the final mosaics in apertures with a radius of 0.3 (F770W), 0.4 (F1000W) and 0.6 arcsec (F1500W, F1800W, F2100W).
+We find non-detections in all analysed images for COSMOS_84213 (F770W26.4, F1800W23.4, at 2), COSMOS_118438 (F770W26.2, F1800W23.5), COSMOS_107923 (F770W26.3, F1000W24.8).
+CEERS_17384 is undetected in the F1000W (25.3), F1500W (24.5), and F2100W (23) bands, but shows a 3 detection in the F770W one (26.6 0.3 AB), which we consider as tentative, because of potential contamination by two bright sources in its vicinity.
+Finally, we used the DAWN JWST Archive (DJA)222https://dawn-cph.github.io/dja/index.html to check whether candidates have imaging observations in NIRCam bands that were not included in M24.
+We find that candidate JADES-GN_9538 is covered by NIRCam medium-band observations in the F182M, F210M and F335M bands.
+It has a marginal (2) detection in the F182M band, it is not detected in the F210M one, and a 6 detection in F335M, resulting in a flat F335M-F356W color.
+We will discuss in Sect.
+4.1 the implication of MIRI and NIRCam medium-bands observations on the redshift probability distribution of our candidates.
+We estimated the MUV of the five candidates by converting to the rest-frame the observed F277W magnitude assuming a redshift z=18 where our redshift selection function peaks, and their UV slope by fitting the F277W, F356W and F444W bands.
+We find two candidates brighter than M-21, with the brightest object in the sample COSMOS_107923 having M-22.7.
+Three of the objects have a red -2, the remaining ones being consistent with a flat or moderately blue UV slope ().
+We have measured their half-light radius by fitting the light distribution in the F277W band with GALIGHT (Ding et al.
+2020; Birrer et al.
+2021) assuming a Sersic (1968) profile with free index and fixing the redshift at the best-fit solution at .
+We find half-light radii consistent with those of observed in galaxies at z10-15 (Westcott et al.
+2025; Ono et al.
+2025).
+Our candidates have a typical half-light radius of 0.2-0.3 kpc.
 We compared our samples with 15 samples selected by other groups in the fields analysed here.
-Our two F200W-dropout candidates in CEERS are not included in the sample of very red sources by Gandolfi et al.
+Our F200W-dropout candidate in CEERS is not included in the sample of very red sources by Gandolfi et al.
 (2025), while 4 of their 5 objects potentially at z15 have indeed colours compatible with our selection window, but have not been included here because of SNR10 in the detection band, and in one case (their A-22691) a marginal (SNR2.2) detection in the ACS stacked image.
@@
 Five of the sources selected by Pérez-González et al.
-(2025) in the MIDIS+NGDEEP observations, including the z19.6 candidate MDS025593 by Pérez-González et al.
+(2025b) in the MIDIS+NGDEEP observations, including the z19.6 candidate MDS025593 by Pérez-González et al.
 (2023b), are not covered by the first-epoch NGDEEP imaging used by M24, the remaining ones being non-detected except their MIDIS-z17-7 (ID=8412 in M24) which is at SNR4, i.e.
 well below our SNR=10 threshold.
-Finally, the 15 sources selected by Hainline et al.
+The 15 sources selected by Hainline et al.
 (2024) in the JADES fields with a counterpart in M24 do not fall within our colour selection window, and are detected at SNR10 except their JADES-GS-53.12692-27.79102 (ID=51718 in M24).
-| ID | R.A. | Dec | F356W | M1500 | Re | zhigh | zlow | |
-| deg. | deg. | AB | kpc | |||||
-| CEERS_15937 | 214.944272 | 52.835847 | 27.15 0.09 | -21.14 | 0.263 0.046 | -1.59 0.35 | 17.2 | 4.6 |
-| CEERS_17384 | 214.853243 | 52.773682 | 27.11 0.08 | -21.06 | 0.317 0.05 | -1.50 0.33 | 17.2 | 4.7 |
-| JADES-GN_9538 | 189.191052 | 62.17421 | 28.90 0.13 | -19.49 | 0.227 0.043 | -1.63 0.40 | 16.1 | 2.8 |
-| COSMOS_31168 | 150.180936 | 2.260756 | 27.67 0.16 | -20.70 | 0.519 0.124 | -2.05 0.39 | 16.2 | 4.0 |
-| COSMOS_35731 | 150.133571 | 2.271020 | 28.54 0.09 | -20.05 | 0.134 0.025 | -1.92 0.17 | 15.9 | 2.8 |
-| COSMOS_76919 | 150.184553 | 2.353510 | 28.71 0.18 | -19.74 | 0.289 0.065 | -2.32 0.44 | 16.1 | 4.0 |
-| COSMOS_84213 | 150.167211 | 2.368995 | 28.14 0.18 | -20.35 | 0.199 0.053 | -2.03 0.40 | 15.6 | 4.0 |
-| COSMOS_107923 | 150.107337 | 2.428780 | 25.48 0.10 | -22.68 | 0.219 0.024 | -1.22 0.26 | 16.3 | 3.3 |
-| COSMOS_118438 | 150.196541 | 2.464192 | 27.86 0.19 | -20.61 | 0.256 0.077 | -2.27 0.47 | 17.1 | 5.0 |
+Finally, the z15 candidate JADES-GN+189.16733+62.31026 by Whitler et al.
+(2025) (ID=56718 in M24) is not selected due to its , i.e.
+slightly below the threshold to enter our ?extended?
+sample presented in Sect.
+A.
+| ID | R.A. | Dec | F356W | M1500 | Re | zhigh | zlow | |||
+| deg. | deg. | AB | kpc | |||||||
+| CEERS_17384 | 214.853243 | 52.77368 | 27.11 0.08 | -21.06 | 0.317 0.05 | -1.50 0.33 | 17.2 | 4.7 | 0.36 | 0.90 |
+| JADES-GN_9538 | 189.191052 | 62.17421 | 28.90 0.13 | -19.49 | 0.227 0.043 | -1.63 0.40 | 16.1 | 2.8 | 0.62 | 0.64 |
+| COSMOS_84213 | 150.167211 | 2.368995 | 28.14 0.18 | -20.35 | 0.199 0.053 | -2.03 0.40 | 15.6 | 4.0 | 2.31 | 1.73 |
+| COSMOS_107923 | 150.107337 | 2.428780 | 25.48 0.10 | -22.68 | 0.219 0.024 | -1.22 0.26 | 16.3 | 3.3 | 0.59 | 0.65 |
+| COSMOS_118438 | 150.196541 | 2.464192 | 27.86 0.19 | -20.61 | 0.256 0.077 | -2.27 0.47 | 17.1 | 5.0 | 1.37 | 0.37 |
 a) ID, coordinates and F356W magnitudes from M24.
@@
 The UV slope is obtained by fitting the F277W, F356W and F444W bands.
-The last two columns show the best-fit solutions obtained with zphot at () and ().
+The last four columns show the best-fit solutions obtained with zphot at () and (), and the relevant values.
 4 A closer scrutiny of the selected candidates
 In this section we analyse in more detail the selected candidates to assess their reliability and evaluate the possibility that their peculiar colours are instead indicative of rare lower redshift interloper populations.
+Although we focus here on the 5 candidates, we remark that very similar arguments can be made for the 5 sources that have been confirmed to be interlopers, as discussed in Sect.
+4.3.
 4.1 The photometric redshift probability distribution
@@
 To summarise, from the combination of codes and assumptions adopted we have obtained eight different for all our candidates.
-The of the nine objects in our reference sample of F200W dropout candidates are shown in the upper panels of Fig.
+The of the five objects in our reference sample of F200W dropout candidates are shown in the upper panels of Fig.
 3.
@@
 We remark that, given the resulting , none of our candidates would pass a selection criterion such as 4 between different redshift solutions that has been often adopted to build luminosity functions at (e.g., Finkelstein et al.
-2023; Harikane et al.
+2024a; Harikane et al.
 2023).
@@
 4).
-We finally note that also depends crucially on the details of the adopted photometry.
+We note that also depends crucially on the details of the adopted photometry.
 Clearly, but somewhat counterintuitively, the additional information provided by the HST photometry decreases the constraining power of .
 Because of the faintness of our candidates, in fact, the HST upper limits are easily satisfied by a wide class of solutions at low redshifts, eventually adding little contribution to the global while increasing the number of degrees of freedom and hence increasing the probability at low redshifts.
+These results are based on the ASTRODEEP-JWST catalogs, which provide homogeneous photometric measurements for all sources.
+We find that the MIRI constraints discussed in Sect.
+3.2 do not significantly change the scenario described above.
+The maintain the same double-peaked nature with both high- and low-redshift solutions when including the MIRI upper limits, and, in the case of CEERS_17384, the tentative F770W detection.
+In fact, the MIRI non-detections are poorly constraining for all our sources, with the exception of COSMOS_107923 for which the limit in the F770W increases the probability of passive solutions with respect to dusty star-forming ones in the low-redshift peak.
+In fact, the MIRI imaging should be 1-2 mag deeper in the F770W and F1000W bands, and 3-4 mag deeper at 15m to yield meaningful constraints for the other candidates.
+Finally, when including additional NIRCam medium-band photometry for JADES-GN_9538 we find an overall similar but with a slightly increased probability at z13-15 due to the marginal detection in the F182M band.
 We take from these results three main lessons.
@@
 Adopting for simplicity the results of the zphot code, we have inspected the physical properties corresponding to the models that populate the low redshift peaks in the .
-The best-fit SEDs at of the nine F200W dropout candidates are shown in Fig.
+The best-fit SEDs at of the F200W dropout candidates are shown in Fig.
 3.
-In the case of the two candidates in CEERS, the low-redshift best fit solution is a ?quiescent?
+In the case of the candidate in CEERS, the low-redshift best fit solution is a ?quiescent?
 galaxy with low specific star–formation rate and large stellar age.
@@
 5.
-We compared the solutions that provide an acceptable fit with probability for the nine F200W dropout candidates to the locus of objects in the same redshift range from the JADES-GS field.
+We compared the solutions that provide an acceptable fit with probability for the F200W-dropout candidates to the locus of objects in the same redshift range from the JADES-GS field.
 The statistically acceptable models populate regions that have a small overlap with the ones occupied by the bulk of sources in the same redshift range.
@@
 2024; Weibel et al.
-2024; Pérez-González et al.
-2024b).
+2025a; Pérez-González et al.
+2025a).
 The available estimates of the stellar mass function of quiescent galaxies do not extend to these low masses, especially at these redshifts (Santini et al.
@@
 2024a).
-4.3 The confirmed interloper UDS_56824 at =6.56
-One of the ten objects originally selected, the F200W-dropout candidate object UDS_56824 at R.A.=34.454893 deg., Dec=-5.215586 deg (see Fig.
-2), has been recently observed with the NIRSpec PRISM in the framework of the CAPERS program.
-The observation has been carried out by adopting a NRSIRS2 readout pattern, standard 3-shutter “slits”, and a 3-point nodding for a total integration time of 5690 secs.
-The data were reduced with the STScI Calibration Pipeline222https://jwst-pipeline.readthedocs.io/en/latest/index.html version 1.13.4 as described in detail in Arrabal Haro et al.
+4.3 The confirmed interlopers
+Five of the ten objects originally selected as F200W-dropouts have been recently observed with the NIRSpec PRISM in the framework of the CAPERS program, and found to be interlopers at z2-7 (empty symbols in Fig.
+2).
+We describe in the following their physical properties as inferred from the combination of NIRSpec spectroscopy (Fig.
+6) and NIRCam photometry (Sect.
+B).
+These observations confirm that contaminants of our selection criteria are found in the populations presented in Sect.
+4.2, namely dusty star—forming or quiescent galaxies with low stellar mass.
+The spectroscopic observations have been carried out by adopting a NRSIRS2 readout pattern, standard 3-shutter “slits”, and a 3-point nodding.
+The data were reduced with the STScI Calibration Pipeline333https://jwst-pipeline.readthedocs.io/en/latest/index.html version 1.13.4 as described in detail in Arrabal Haro et al.
 (2023a) (see also Castellano et al.
@@
 2025).
-Although the spectrum covers only the region at 3m, the redshift can be accurately measured to be =6.560.01 from the weighted average centroid of the H, H, [O III] lines which are clearly detected (Fig.
-6).
-We measured line fluxes with a Gaussian fit after linearly extrapolating the continuum emission at the line position (see Napolitano et al.
-2024, 2025, for details) finding FHα=16.2 2.1 10, FHβ=2.1 0.9 10, F[OIII]4959=2.1 1.1 10, and F[OIII]5007=1.26 0.18 10.
-The Balmer ratio implies a high dust attenuation with =0.850.46 when assuming a Calzetti et al.
-(2000) attenuation law.
+We determine the spectroscopic redshift from the weighted average centroid of the detected emission lines, and measured line fluxes with a Gaussian fit after linearly extrapolating the continuum emission at the line position (see Napolitano et al.
+2024, 2025, for details).
+4.3.1 COSMOS_31168 at =2.29
+The spectrum of COSMOS_31168 (R.A.= 150.1809331 deg., Dec=2.2607514 deg.) shows one clear emission line at 1.64m, a lower significance emission feature at 2.15m and a faint continuum.
+The most plausible redshift solution is =2.294, interpreting the aforementioned emission lines as the unresolved [O III] doublet (with flux 9.71.9 10) and H (=3.10.5 10).
+The H line is not detected, with a 2 flux upper limit of 0.75 10.
+The corresponding limit on the Balmer ratio results in =0.32.
+We performed an SED-fitting of the object with zphot by fixing the redshift at the spectroscopic value and adopting the aforementioned constraint on , finding that the object is fitted as a low-mass (log = 8), low SFR (=0.07 M⊙ yr-1) dusty object.
+4.3.2 COSMOS_35731 at =3.66
+The CAPERS spectrum of this source (R.A.= 150.133574 deg., Dec=2.2710158 deg.) has prominent H (28.60.2 10), [O III] (8.70.2 10) and H (4.80.2 10) emission, yielding a robust spectroscopic redshift =3.657.
+The Balmer decrement implies a high extinction (0.60.04), with a corrected SFR measured from the H luminosity of 1.9 0.1 M⊙ yr-1.
+The SED-fitting performed at fixed redshift and constraining within the 2 range inferred from spectroscopy yields a stellar mass of only log = 7.6.
+4.3.3 CEERS_15937 at =4.91
+The spectrum of CEERS_15937 (R.A.= 214.9442826 deg., Dec= 52.8358477 deg.) shows continuum emission with a clear break at 2.15m.
+The shape of the continuum and the faint emission detectable down to 1m shows that the object is not at very high-redshift.
+Due to the lack of evident emission features, we fitted the spectrum with BAGPIPES assuming a double power-law SFH, and leaving the redshift as a free parameter.
+We find that the object is fitted by a low-mass (log = 8.9 0.05) galaxy at =4.91, with low extinction (0.02) and negligible SFR, corresponding to an upper limit of sSFR-11.4 Gyr-1.
+4.3.4 COSMOS_76919 at =5.18
+This object (R.A.= 150.1845568 deg., Dec= 2.3535079 deg.) is constrained to be at =5.182 from the detection of the H (2.20.5 10), [O III] (4.80.3 10), and H (0.4 0.2 10) lines.
+Similarly to COSMOS_31168 and COSMOS_35731, the object turns out to be a dusty (0.50.3), low-mass (log = 8.2 0.4), and low SFR (=0.3 0.1 M⊙ yr-1) object on the basis of the Balmer decrement, H luminosity and constrined SED-fitting performed with zphot.
+4.3.5 UDS_56824 at =6.56
+The spectrum of object UDS_56824 (R.A.=34.454893 deg., Dec=-5.215586 deg) covers only the region at 3m, with a total integration time of 5690 secs.
+The redshift can be accurately measured to be =6.560.01 from the H, H, [O III] lines, which are clearly detected with fluxes FHα=16.2 2.1 10, FHβ=2.1 0.9 10, F[OIII]4959=2.1 1.1 10, and F[OIII]5007=1.26 0.18 10.
+The Balmer ratio implies a high dust attenuation with =0.850.46.
 Its [O III]/H=6.02.7 puts the object in the star-forming region of the mass-excitation diagram (Juneau et al.
@@
 6).
-This galaxy is fainter than all SCUBA-2-selected galaxies in UDS (Geach et al.
-2017) and all previous ALMA detections in the field reported in Dudzevičiūtė et al.
-(2020), suggesting that a significant population of dusty galaxies at z6 may have been missed by previous submillimeter surveys.
 The measured ALMA flux density was then combined with NIRCam photometry to perform an energy-balance SED fitting using BAGPIPES.
@@
 Nevertheless, even in this scenario, this galaxy would stand out due to its high dust attenuation compared to UV-selected galaxies.
-The combination of spectroscopic and photometric information demonstrates that UDS_56824 is an example of a highly attenuated high-redshift starburst galaxy.
-Its large UV attenuation 4.5 mag, makes it a significant outlier in the relation (McLure et al.
-2018), similarly to CEERS-93316 (Arrabal Haro et al.
-2023b) and CEERS-14821 (Bisigello et al.
+4.4 Critical assessment
+These findings are consistent with our analysis in Sect.
+4.2 indicating dusty star-forming galaxies and low-mass passive objects up to z7-8 as a potential source of contamination of the F200W-dropout sample.
+However, it is apparent that a proper characterization of these interloper populations is extremely difficult and requires spectroscopic follow-up observations.
+In fact, CEERS_15937 is the only case for which the yielded an alternative solution in good agreement with the spectroscopic value (Fig.
+16).
+Instead, the redshifts of the other four objects do not match the primary, low-redshift peaks in our , although the most likely alternative solutions are similarly dusty, star-forming objects whose broad-band fluxes are boosted by different combinations of emission lines.
+Similarly to CEERS-93316 (Arrabal Haro et al.
+2023b), the SEDs of COSMOS_31168, COSMOS_35731, COSMOS_76919 and UDS_56824 resemble that of z15 LBGs due to the combination of a red, attenuated continuum and line emission.
+Instead, CEERS_15937 enters the F200W-droput selection due to the well-known ambiguity between Lyman and Balmer break detection through broad-band photometry.
+Their position in the vs.
+, and in the vs.
+planes are in line with the expectations based on the (Fig.
+5).
+The four dusty star-forming interlopers are significant outliers in the relation (McLure et al.
+2018), similarly to CEERS-93316 and CEERS-14821 (Bisigello et al.
 2025).
+Object UDS_56824 appears to be an extreme case in terms of redshift and stellar mass compared to the locus of alternative solutions of the 15 candidates.
+The other known interlopers at z2-5 are similar to UDS_56824 in terms of extinction and sSFR, but with a lower stellar mass log9.
 In fact, UDS_56824 shows that even at z6.5, some galaxies may be dominated by dust-obscured star formation (note that the uncorrected H-based SFR is 2-4 M⊙ yr-1).
-The other known interloper CEERS-93316 at z=4.9 is remarkably similar to UDS_56824 in terms of both extinction (0.6 mag) and SFR (30 M⊙ yr-1), but with a lower stellar mass log8.4-9 (Arrabal Haro et al.
-2023b).
-These findings are consistent with our analysis in Sect.
-4.2 pointing toward dusty star-forming galaxies up to z7-8 as a potential source of contamination of the F200W-dropout sample.
-Similarly to CEERS-93316, the SED of UDS_56824 resembles that of a z15 LBG due to the combination of a red, attenuated continuum and line emission.
-However, it is apparent that a proper characterization of these interloper populations is extremely difficult and requires spectroscopic follow-up observations.
-In fact, a redshift of z6.5 is a secondary peak in our (Fig.
-7) but the most likely alternative solution was a similarly dusty, star-forming object with log9 at z4 whose broad-band fluxes are boosted by a different combination of emission lines.
-Moreover, the properties of UDS_56824 appear extreme compared to the locus of alternative solutions of the other 15 candidates, while CEERS-93316 is more in line with the expectations (Fig.
-5).
-4.4 Critical assessment of the photometric sample
-We exploit the analysis described above to perform a detailed evaluation of the objects in our reference sample of F200W-dropouts.
-We first note that four objects in our sample consistently show a probability for the high-redshift solution which is significantly lower than any of the low/intermediate redshift peaks: COSMOS_31168, COSMOS_76919, COSMOS_84213 and COSMOS_118438.
-In the case of COSMOS_31168 and COSMOS_84213, this is likely due to a marginal detection (SNR2) in one HST band each, albeit both objects are non-detected in both the ACS and WFC3 stacks that we used for our selection.
-In addition, both these sources as well as COSMOS_76919 show a drop in the F410M band similarly to UDS_56824, thus making dusty star-forming solutions more likely.
-On the contrary, COSMOS_118438 has a slightly higher flux in F410M than in F356W and F444W, which also leads SED-fitting codes to prefer strong-line emitting templates at z6.
-In fact, we tested that the probability of the solution increases after removing the F410M from the fit for all the aforementioned sources.
-The remaining five sources in our sample have high-redshift solutions as significant as the low-redshift ones in most of the SED-fitting runs.
-In particular, CEERS_15937, CEERS_17384 and COSMOS_107923 show a high consistency among all codes and recipes.
-However, there are compelling reasons to consider their reliability with caution.
-The two CEERS objects appear somewhat similar to CEERS-93316, as they have a clear peak at and a red continuum in the F277W/F356W/F444W bands, leading us to suspect that they are members of the same 5 overdensity (Naidu et al.
+Finally, this ALMA-detected galaxy is fainter than all SCUBA-2-selected sources in UDS (Geach et al.
+2017) and all previous ALMA detections in the field reported in Dudzevičiūtė et al.
+(2020), suggesting that a significant population of dusty galaxies at z6 may have been missed by previous submillimeter surveys.
+Object CEERS_15937 at z=4.91 is remarkable, being among the most distant confirmed passive objects.
+Most importantly, it is the least massive known to date, the only two comparable cases being GS-z5-Q1 and COS-z5-Q1 with a 3 times larger stellar mass (log9.5-9.6, Baker et al.
+2025).
+Similarly to GS-z5-Q1 and COS-z5-Q1, CEERS_15937 is a member of a known 5 overdensity (Naidu et al.
 2022b; Arrabal Haro et al.
@@
 2025) galaxies.
+The number density and formation pathway of low-mass quiescent galaxies at high-redshift are presently unknown (e.g., Merlin et al.
+2025).
+An in-depth investigation of CEERS_15937 and similar objects may shed light on the role of enviroment at early times and on their connection with the recently discovered class of ?mini-quenched?
+galaxies at high-redshift (Strait et al.
+2023; Looser et al.
+2024).
+We exploit the analysis described above to perform a detailed evaluation of the objects in our reference sample of F200W-dropouts.
+We first note that two objects in our sample consistently show a probability for the high-redshift solution which is significantly lower than any of the low/intermediate redshift peaks: COSMOS_84213 and COSMOS_118438.
+In the case of COSMOS_84213, this is likely due to a marginal detection (SNR2) in one HST band each, albeit it is non-detected in both the ACS and WFC3 stacks that we used for our selection.
+In addition, this source show a drop in the F410M band similarly to COSMOS_31168, COSMOS_76919, and UDS_56824, thus making dusty star-forming solutions more likely.
+On the contrary, COSMOS_118438 has a slightly higher flux in F410M than in F356W and F444W, which also leads SED-fitting codes to prefer strong-line emitting templates at z6.
+In fact, we tested that the probability of the solution increases after removing the F410M from the fit for all the aforementioned sources.
+The remaining three sources in our sample have high-redshift solutions as significant as the low-redshift ones in most of the SED-fitting runs.
+In particular, CEERS_17384 and COSMOS_107923 show a high consistency among all codes and recipes.
+However, there are compelling reasons to consider their reliability with caution.
+The CEERS object appear somewhat similar to CEERS_15937, as it has a clear peak corresponding to a passive solution at , leading us to suspect that it is a member of the same 5 overdensity (Naidu et al.
+2022b; Arrabal Haro et al.
+2023b).
 Finally, COSMOS_107923, albeit being the brightest in the sample falls in a region of relatively shallow image depth and therefore has a corresponding limited sampling of the spectral break.
 This source also has -1.2 which is redder than the UV slopes measured in spectroscopically confirmed objects at (Roberts-Borsani et al.
-2024b, a), with the only notable exception of the X-ray emitting AGN GHZ9 at z=10.145 (=-1.10.12, Napolitano et al.
+2024, 2025a), with the only notable exception of the X-ray emitting AGN GHZ9 at z=10.145 (=-1.10.12, Napolitano et al.
 2024).
@@
 In order to take into account the caveats discussed above on the potential contamination from rare classes of interlopers, we consider two opposite scenarios regarding the reliability of our F200W-dropout sample.
-In Case 1, we assume that all our 9 candidates are at , as indicated by the selection function of Fig.
+In Case 1, we assume that all our five candidates are at , as indicated by the selection function of Fig.
 1.
@@
 However, for the reasons described above, we believe that the knowledge of the population of potential interlopers in LBG selections at these extreme redshifts is too uncertain at the moment, and we prefer to bracket the various options with the two opposite scenarios described above.
-The results are reported in Tables 2 (z=15-20) and 3 (z=20-30), and shown in Fig. 8 and 9, respectively.
-The comparison to available measurements of the LF at clearly shows that if all our 9 F200W-dropout candidates are at (Case 1), the LF continues a trend of very slow evolution, similar to recent findings at (Harikane et al.
+The results are reported in Tables 2 (z=15-20) and 3 (z=20-30), and shown in Fig. 7 and 8, respectively.
+The comparison to available measurements of the LF at clearly shows that if all our 5 F200W-dropout candidates are at (Case 1), the LF continues a trend of very slow evolution, similar to recent findings at (Harikane et al.
 2025; Donnan et al.
@@
 2024; Finkelstein et al.
-2024).
+2024b).
 Our estimates would be in agreement with the measurements at z=16-18 by Harikane et al.
@@
 (2025), Kokorev et al.
-(2024) and Pérez-González et al.
-(2025) in the same luminosity range.
+(2025) Pérez-González et al.
+(2025b), and Weibel et al.
+(2025b) in the same luminosity range.
 Most importantly, a similar scenario remains valid if we consider partial, but not complete, contamination of our sample and/or successful confirmation of some candidates in our extended sample.
 In fact, every point in the LF is originated by a small number of observed galaxies falling in that luminosity bin (typically 1 or 2), such that even if a small fraction of candidates is indeed at , the corresponding density in the considered bin would be comparable to the available estimates at 14.
-In addition, three of our candidates are brighter than the highest-redshift secure galaxy JADES-GS-z14-0 (MUV=-20.81, Carniani et al.
-2024a): if confirmed, they would imply a number density at M-21 higher than current measurements at 14.
+In addition, two of our candidates are brighter than the highest-redshift secure galaxy JADES-GS-z14-0 (MUV=-20.81, Carniani et al.
+2025): if confirmed, they would imply a number density at M-21 higher than current measurements at 14.
 Needless to say, if we assume that none of our candidates are genuinely at (Case 2), the upper limits simply indicate that the actual LF is located at lower densities and luminosities, to an extent that we cannot establish with the existing data.
@@
 The derived upper limits are consistent with the z25 estimates by Pérez-González et al.
-(2025), and somewhat above the existing measurements at and at , due to the reduced effective volume sampled by our surveys.
+(2025b), and somewhat above the existing measurements at and at , due to the reduced effective volume sampled by our surveys.
 This result demonstrates that significantly wider/deeper areas are in principle necessary to sample this redshift range - a point that will be further expanded below.
@@
 | -22.5 | 1 | 0.06 | 0.12 |
-| -21.5 | 2 | 0.14 | 0.13 |
-| -20.5 | 4 | 0.50 | 0.23 |
-| -19.5 | 2 | 1.2 | 1.1 |
+| -21.5 | 1 | 0.07 | 0.13 |
+| -20.5 | 2 | 0.25 | 0.23 |
+| -19.5 | 1 | 0.6 | 1.1 |
 | -18.5 | 0 | 5.4 | 5.4 |
@@
 We let then evolve up to the LF at under the assumption of a non-evolving , which is shown in the same panel of Fig.
-10.
+9.
 In practice, the entire evolution of the LF is driven by the corresponding evolution of the Press & Schechter HMF.
@@
 Notably, the resulting UV LFs at 15 fall below our measurements (Fig.
-10, right panel).
+9, right panel).
 To further include effects that maximize the UV luminosity associated with a given dark matter halo, we also allowed for a stochastic fluctuation of the star formation rate, adopting the simplified description proposed by Kravtsov & Belokurov (2024).
@@
 Only the additional effect of extreme stochasticity, highlighted by a shaded region in Fig.
-10, allows this simple empirical model to match the number density inferred from our reference F200W-dropout sample (consistently with the results by Pallottini & Ferrara 2023).
+9, allows this simple empirical model to match the number density inferred from our reference F200W-dropout sample (consistently with the results by Pallottini & Ferrara 2023).
 A similar result may be obtained by allowing for top-heavy, or flat IMFs, which can increase by up to a factor of 10 compared to a Chabrier IMF.
@@
 2023; Li et al.
-2024) the excess of bright galaxies is explained as the result of high densities and low metallicities yielding a extremly high star-formation efficiency at cosmic dawn.
+2024) the excess of bright galaxies is explained as the result of high densities and low metallicities yielding a extremly high star-formation efficiency at cosmic dawn444The predictions used in the present work incorporate the improved estimate of the UV luminosity based on the median SFH, as described in Dekel et al.
+(2025)..
 The DELPHI semi-analytic model (SAM) based on cold gas fractions and star formation efficiencies sampled from the sphinx simulations (Mauerhofer et al.
@@
 2024).
-Finally, the recently updated Santa Cruz semi-analytic model (Somerville et al.
-in prep., see also Yung et al.
-2024a) was run on dark matter halo merger trees extracted from the GUREFT simulations (Yung et al.
+Finally, the recently updated Santa Cruz semi-analytic model (Yung et al.
+2024a; Somerville et al.
+2025) was run on dark matter halo merger trees extracted from the GUREFT simulations (Yung et al.
 2024b), and incorporates a star formation efficiency that increases with increasing gas surface density, motivated by results from molecular cloud-scale simulations with radiative transfer.
@@
 Consistently with the simple empirical predictions described above, all these theoretical scenarios point towards a strong evolution at (Fig.
-11).
-Both the AFM model, and the FFB one with star-formation efficiency , provide a good match to the z12 UV LF but predict an abundance of galaxies at z15 lower than our Case 1 LF.
-Similarly, the CAT SAM based on a standard IMF (Trinca et al.
+10).
+Both the AFM model, and the FFB one with star-formation efficiency , provide a good match to the z12 UV LF but while the AFM predicts an abundance of galaxies at z15 lower than our Case 1 LF, the FFB model with is consistent with our estimates at M-22.
+The CAT SAM based on a standard IMF (Trinca et al.
 2024), the DELPHI ?fiducial?
 model, and the Santa Cruz SAM with =0.1, fall below our estimates.
-Interestingly, all models are inconsistent at 2 with the Case 1 abundance at M-20, but the tension is alleviated when assuming a change in physical properties at z15.
+Interestingly, 2 tensions are found at the brightest magnitudes, i.e.
+at M-22 in the case of the FFB model, and at M-20 for the others.
+The tensions are apparently alleviated when assuming a change in physical properties.
 The DELPHI eIMF and eSFE models, and the CAT model assuming an evolving IMF, are partially consistent with our estimates at -18.5–19.5.
-An evolution of the dense gas fraction from 0.1 at z=13 to 0.5 could match our estimates at all luminosities according to the Santa Cruz SAM.
-The FFB model with is also consistent with our Case 1 results at -20, but lower SFEs may also be viable if galaxies at 15 follow a relation between stellar mass and UV luminosity different than the z10 one adopted by Li et al.
-(2024).
+An evolution of the dense gas fraction from 0.1 at z=13 to 0.5 could match our estimates at all luminosities according to the Santa Cruz SAM, while the FFB model with would match the observed Case 1 abundance in the brightest bin.
 These comparisons help us to put the results presented in this paper in context, leading to our main conclusions:
@@
 1) with a contamination rate of 0.5% of the considered templates, with a significant gain in the accessible color space.
+Our analysis is in good agreement with the results by Adams et al.
+(2025), showing that doubling the exposure time in the bands sampling below the break is the most efficient way to reduce contamination in a standard high-redshift survey.
 We then addressed the question of the survey area required to detect a given number of sources as predicted by the various theoretical models discussed in the previous section.
 The results are shown in Fig.
-12.
+11.
 Finding at least one object at the very bright end (M-20) of the UV LF at requires reaching a continuum magnitude around 28 (at SNR=10) over an area of sq.arcmin in most of the scenarios.
@@
 On the basis of mock observations mimicking the properties of our dataset, we have designed specific renditions of the Lyman-break selection technique that efficiently identify galaxy candidates in the redshift ranges and .
-We isolated nine candidates at , while no objects are found at .
+We isolated five candidates at , while no objects are found at .
 A closer inspection of the selected candidates shows that despite exhibiting a 1.5 mag break, the selected objects consistently display multimodal redshift probability distributions across different SED-fitting codes and methodologies.
@@
 Most importantly, they correspond to populations of low-mass () quiescent or dusty galaxies that have not been thoroughly investigated so far.
-This result is corroborated by the spectral properties of object UDS_56824, a confirmed interloper of our F200W-dropout selection, which is found to be a dusty () starburst galaxy at with mass .
+This result is corroborated by the spectral properties of five confirmed interlopers of our F200W-dropout selection.
+Four objects are found to be dusty star-forming galaxies, including three low-mass (7.6-8.8) objects at z2-5, and a highly attenuated () starburst galaxy at with mass .
+The fifth confirmed interloper, CEERS_15937 at z=4.91, is found to be the first spectroscopically confirmed passive galaxy with log 9 at high-redshift.
 These results imply that while our candidates are, in principle, credible objects at , none of them would pass stringent selection criteria based on between different redshift solutions.
@@
 In particular, the confirmation of bright (M-21) candidates would require deep revisions of our theoretical framework, and might be even in contrast with any plausible model under a standard –CDM cosmological scenario.
-However, the tension with theoretical models is so significant that even if only a small fraction of the candidates is confirmed to be at , a further evolution of physical properties, such as IMF or star-formation efficiency would be required to explain the observed number densities.
+The tension with theoretical models implies that even if only a small fraction of the candidates is confirmed to be at , a further evolution of physical properties, such as IMF or star-formation efficiency would be required to explain the observed number densities.
 If instead all the analysed candidates are interlopers, we are forced to conclude that future surveys will need to cover much wider areas to secure the selection of bright galaxies significantly beyond .
@@
 Acknowledgements.
+We thank the referee for the constructive comments that helped us improve the manuscript.
 We thank J.
@@
 We acknowledge financial support from NASA through grant JWST-ERS-1324.
-Support was also provided by the PRIN 2022 MUR project 2022CB3PJ3 – First Light And Galaxy aSsembly (FLAGS) funded by the European Union – Next Generation EU, and by INAF Mini-grant “Reionization and Fundamental Cosmology with High-Redshift Galaxies.” P.G.P.-G.
+Support was also provided by the PRIN 2022 MUR project 2022CB3PJ3 – First Light And Galaxy aSsembly (FLAGS) funded by the European Union – Next Generation EU, by INAF Mini-grant “Reionization and Fundamental Cosmology with High-Redshift Galaxies”, and by INAF GO Grant ”Revealing the nature of bright galaxies at cosmic dawn with deep JWST spectroscopy.” P.G.P.-G.
 and L.C.
@@
 acknowledges support from grant “Progetti per Avvio alla Ricerca - Tipo 1, Unveiling Cosmic Dawn: Galaxy Evolution with CAPERS” (AR1241906F947685).
-RA acknowledges financial support from grants PID2023-147386NB-I00 ”XTREM” and PID2022-136598NB-C32 “Estallidos8”, funded by MCIN/AEI/10.13039/501100011033 and ERDF/EU, and the State Agency for Research of the Spanish MCIU through ‘Center of Excellence Severo Ochoa’ award to the IAA-CSIC (SEV-2017-0709) and CEX2021-001131-S.
+RA acknowledges financial support from grants PID2023-147386NB-I00 ”XTREM” and PID2022-136598NB-C32 “Estallidos8”, funded by MCIU/AEI/10.13039/501100011033 and ERDF/EU, and the State Agency for Research of the Spanish MCIU through ‘Center of Excellence Severo Ochoa’ award to the IAA-CSIC (SEV-2017-0709) and CEX2021-001131-S.
 F.P.
@@
 The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST.
-These observations are associated with programs #1324, 1345, 1180, 1210, 1837, 2561, 2756, 3990, and 6368.References
+These observations are associated with programs #1324, 1345, 1180, 1210, 1837, 2561, 2756, 3794, 3990, 5893, and 6368.
+This paper makes use of the following ALMA data: ADS/JAO.ALMA#2015.1.01074.S.
+ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile.
+The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ.
+Some of the data products presented herein were retrieved from the Dawn JWST Archive (DJA).
+DJA is an initiative of the Cosmic Dawn Center (DAWN), which is funded by the Danish National Research Foundation under grant DNRF140.References
+- Adams et al. (2025) Adams, N. J., Austin, D., Harvey, T., et al. 2025, MNRAS, 542, 1705
 - Algera et al. (2024) Algera, H. S. B., Inami, H., De Looze, I., et al. 2024, MNRAS, 533, 3098
@@
 - Bagley et al. (2024) Bagley, M. B., Pirzkal, N., Finkelstein, S. L., et al. 2024, ApJ, 965, L6
-- Bergamini et al. (2023) Bergamini, P., Acebron, A., Grillo, C., et al. 2023, arXiv e-prints, arXiv:2303.10210
+- Baker et al. (2025) Baker, W. M., Ito, K., Valentino, F., et al. 2025, arXiv e-prints, arXiv:2509.09761
+- Bergamini et al. (2023) Bergamini, P., Acebron, A., Grillo, C., et al. 2023, ApJ, 952, 84
 - Bertin & Arnouts (1996) Bertin, E. & Arnouts, S. 1996, A&AS, 117, 393
-- Bezanson et al. (2022) Bezanson, R., Labbe, I., Whitaker, K. E., et al. 2022, arXiv e-prints, arXiv:2212.04026
+- Bezanson et al. (2024) Bezanson, R., Labbe, I., Whitaker, K. E., et al. 2024, ApJ, 974, 92
 - Birrer et al. (2021) Birrer, S., Shajib, A., Gilman, D., et al. 2021, The Journal of Open Source Software, 6, 3283
@@
 - Carnall et al. (2018) Carnall, A. C., McLure, R. J., Dunlop, J. S., & Davé, R. 2018, MNRAS, 480, 4379
-- Carniani et al. (2024a) Carniani, S., D’Eugenio, F., Ji, X., et al. 2024a, arXiv e-prints, arXiv:2409.20533
-- Carniani et al. (2024b) Carniani, S., Hainline, K., D’Eugenio, F., et al. 2024b, Nature, 633, 318
+- Carniani et al. (2025) Carniani, S., D’Eugenio, F., Ji, X., et al. 2025, A&A, 696, A87
+- Carniani et al. (2024) Carniani, S., Hainline, K., D’Eugenio, F., et al. 2024, Nature, 633, 318
 - Casey et al. (2024) Casey, C. M., Akins, H. B., Shuntov, M., et al. 2024, ApJ, 965, 98
@@
 - Cole et al. (2025) Cole, J. W., Papovich, C., Finkelstein, S. L., et al. 2025, ApJ, 979, 193
-- Conselice et al. (2024) Conselice, C. J., Adams, N., Harvey, T., et al. 2024, arXiv e-prints, arXiv:2407.14973
+- Conselice et al. (2025) Conselice, C. J., Adams, N., Harvey, T., et al. 2025, ApJ, 983, 30
 - Correa et al. (2015) Correa, C. A., Wyithe, J. S. B., Schaye, J., & Duffy, A. R. 2015, MNRAS, 450, 1514
@@
 - de Graaff et al. (2025) de Graaff, A., Setton, D. J., Brammer, G., et al. 2025, Nature Astronomy, 9, 280
+- Dekel et al. (2025) Dekel, A., Mandelker, N., Li, Z., et al. 2025, arXiv e-prints, arXiv:2506.11664
 - Dekel et al. (2023) Dekel, A., Sarkar, K. C., Birnboim, Y., Mandelker, N., & Li, Z. 2023, MNRAS, 523, 3201
@@
 - Ferrara et al. (2025b) Ferrara, A., Pallottini, A., & Sommovigo, L. 2025b, A&A, 694, A286
-- Finkelstein et al. (2025) Finkelstein, S. L., Bagley, M. B., Arrabal Haro, P., et al. 2025, arXiv e-prints, arXiv:2501.04085
+- Finkelstein et al. (2025) Finkelstein, S. L., Bagley, M. B., Arrabal Haro, P., et al. 2025, ApJ, 983, L4
 - Finkelstein et al. (2022) Finkelstein, S. L., Bagley, M. B., Arrabal Haro, P., et al. 2022, ApJ, 940, L55
-- Finkelstein et al. (2023) Finkelstein, S. L., Leung, G. C. K., Bagley, M. B., et al. 2023, arXiv e-prints, arXiv:2311.04279
-- Finkelstein et al. (2024) Finkelstein, S. L., Leung, G. C. K., Bagley, M. B., et al. 2024, ApJ, 969, L2
+- Finkelstein et al. (2024a) Finkelstein, S. L., Leung, G. C. K., Bagley, M. B., et al. 2024a, ApJ, 969, L2
+- Finkelstein et al. (2024b) Finkelstein, S. L., Leung, G. C. K., Bagley, M. B., et al. 2024b, ApJ, 969, L2
 - Fiore et al. (2023) Fiore, F., Ferrara, A., Bischetti, M., Feruglio, C., & Travascio, A. 2023, ApJ, 943, L27
@@
 - Harikane et al. (2023) Harikane, Y., Ouchi, M., Oguri, M., et al. 2023, ApJS, 265, 5
+- Haslbauer et al. (2022) Haslbauer, M., Kroupa, P., Zonoozi, A. H., & Haghi, H. 2022, ApJ, 939, L31
 - Holwerda et al. (2018) Holwerda, B. W., Bridge, J. S., Ryan, R., et al. 2018, A&A, 620, A132
@@
 - Koekemoer et al. (2011) Koekemoer, A. M., Faber, S. M., Ferguson, H. C., et al. 2011, ApJS, 197, 36
-- Kokorev et al. (2024) Kokorev, V., Atek, H., Chisholm, J., et al. 2024, arXiv e-prints, arXiv:2411.13640
+- Kokorev et al. (2025) Kokorev, V., Atek, H., Chisholm, J., et al. 2025, ApJ, 983, L22
 - Kravtsov & Belokurov (2024) Kravtsov, A. & Belokurov, V. 2024, arXiv e-prints, arXiv:2405.04578
 - Kron (1980) Kron, R. G. 1980, ApJS, 43, 305
-- Larson et al. (2022) Larson, R. L., Hutchison, T. A., Bagley, M., et al. 2022, arXiv e-prints, arXiv:2211.10035
+- Larson et al. (2023) Larson, R. L., Hutchison, T. A., Bagley, M., et al. 2023, ApJ, 958, 141
 - Leung et al. (2023) Leung, G. C. K., Bagley, M. B., Finkelstein, S. L., et al. 2023, ApJ, 954, L46
 - Li et al. (2024) Li, Z., Dekel, A., Sarkar, K. C., et al. 2024, A&A, 690, A108
+- Looser et al. (2024) Looser, T. J., D’Eugenio, F., Maiolino, R., et al. 2024, Nature, 629, 53
 - Marley et al. (2021) Marley, M. S., Saumon, D., Visscher, C., et al. 2021, ApJ, 920, 85
@@
 - Matteri et al. (2025) Matteri, A., Ferrara, A., & Pallottini, A. 2025, arXiv e-prints, arXiv:2503.18850
-- Mauerhofer et al. (2025) Mauerhofer, V., Dayal, P., Haehnelt, M. G., et al. 2025, arXiv e-prints, arXiv:2502.02647
+- Mauerhofer et al. (2025) Mauerhofer, V., Dayal, P., Haehnelt, M. G., et al. 2025, A&A, 696, A157
 - McGaugh et al. (2024) McGaugh, S. S., Schombert, J. M., Lelli, F., & Franck, J. 2024, ApJ, 976, 13
@@
 - Menci et al. (2024) Menci, N., Sen, A. A., & Castellano, M. 2024, ApJ, 976, 227
-- Merlin et al. (2022) Merlin, E., Bonchi, A., Paris, D., et al. 2022, arXiv e-prints, arXiv:2207.11701
+- Merlin et al. (2022) Merlin, E., Bonchi, A., Paris, D., et al. 2022, ApJ, 938, L14
+- Merlin et al. (2025) Merlin, E., Fortuni, F., Calabró, A., et al. 2025, arXiv e-prints, arXiv:2509.09764
 - Merlin et al. (2019) Merlin, E., Pilo, S., Fontana, A., et al. 2019, A&A, 622, A169
@@
 - Mitsuhashi et al. (2024) Mitsuhashi, I., Harikane, Y., Bauer, F. E., et al. 2024, ApJ, 971, 161
-- Morishita et al. (2024) Morishita, T., Mason, C. A., Kreilgaard, K. C., et al. 2024, arXiv e-prints, arXiv:2412.04211
+- Morishita et al. (2025) Morishita, T., Mason, C. A., Kreilgaard, K. C., et al. 2025, ApJ, 983, 152
 - Naidu et al. (2022a) Naidu, R. P., Matthee, J., Oesch, P. A., et al. 2022a, MNRAS, 510, 4582
+- Naidu et al. (2025) Naidu, R. P., Oesch, P. A., Brammer, G., et al. 2025, arXiv e-prints, arXiv:2505.11263
 - Naidu et al. (2022b) Naidu, R. P., Oesch, P. A., Setton, D. J., et al. 2022b, arXiv e-prints, arXiv:2208.02794
@@
 - Oke & Gunn (1983) Oke, J. B. & Gunn, J. E. 1983, ApJ, 266, 713
-- Ono et al. (2023) Ono, Y., Harikane, Y., Ouchi, M., et al. 2023, ApJ, 951, 72
 - Ono et al. (2025) Ono, Y., Ouchi, M., Harikane, Y., et al. 2025, arXiv e-prints, arXiv:2502.08885
@@
 - Paris et al. (2023) Paris, D., Merlin, E., Fontana, A., et al. 2023, ApJ, 952, 20
-- Peng et al. (2010) Peng, C. Y., Ho, L. C., Impey, C. D., & Rix, H.-W. 2010, AJ, 139, 2097
 - Pérez-González et al. (2023a) Pérez-González, P. G., Barro, G., Annunziatella, M., et al. 2023a, ApJ, 946, L16
 - Pérez-González et al. (2024a) Pérez-González, P. G., Barro, G., Rieke, G. H., et al. 2024a, ApJ, 968, 4
-- Pérez-González et al. (2023b) Pérez-González, P. G., Costantin, L., Langeroodi, D., et al. 2023b, arXiv e-prints, arXiv:2302.02429
-- Pérez-González et al. (2024b) Pérez-González, P. G., D‘Eugenio, F., Rodríguez del Pino, B., et al. 2024b, arXiv e-prints, arXiv:2405.03744
-- Pérez-González et al. (2025) Pérez-González, P. G., Östlin, G., Costantin, L., et al. 2025, arXiv e-prints, arXiv:2503.15594
-- Pérez-González et al. (2024c) Pérez-González, P. G., Rinaldi, P., Caputi, K. I., et al. 2024c, ApJ, 969, L10
+- Pérez-González et al. (2023b) Pérez-González, P. G., Costantin, L., Langeroodi, D., et al. 2023b, ApJ, 951, L1
+- Pérez-González et al. (2025a) Pérez-González, P. G., D’Eugenio, F., Rodríguez del Pino, B., et al. 2025a, Nature Astronomy, 9, 1240
+- Pérez-González et al. (2025b) Pérez-González, P. G., Östlin, G., Costantin, L., et al. 2025b, ApJ, 991, 179
+- Pérez-González et al. (2024b) Pérez-González, P. G., Rinaldi, P., Caputi, K. I., et al. 2024b, ApJ, 969, L10
 - Rinaldi et al. (2024) Rinaldi, P., Navarro-Carrera, R., Caputi, K. I., et al. 2024, arXiv e-prints, arXiv:2406.13554
-- Roberts-Borsani et al. (2024a) Roberts-Borsani, G., Bagley, M., Rojas-Ruiz, S., et al. 2024a, arXiv e-prints, arXiv:2407.17551
-- Roberts-Borsani et al. (2024b) Roberts-Borsani, G., Treu, T., Shapley, A., et al. 2024b, arXiv e-prints, arXiv:2403.07103
+- Roberts-Borsani et al. (2025a) Roberts-Borsani, G., Bagley, M., Rojas-Ruiz, S., et al. 2025a, ApJ, 983, 18
+- Roberts-Borsani et al. (2025b) Roberts-Borsani, G., Oesch, P., Ellis, R., et al. 2025b, arXiv e-prints, arXiv:2508.21708
+- Roberts-Borsani et al. (2024) Roberts-Borsani, G., Treu, T., Shapley, A., et al. 2024, ApJ, 976, 193
 - Robertson et al. (2024) Robertson, B., Johnson, B. D., Tacchella, S., et al. 2024, ApJ, 970, 31
@@
 - Sheth & Tormen (1999) Sheth, R. K. & Tormen, G. 1999, MNRAS, 308, 119
+- Somerville et al. (2025) Somerville, R. S., Yung, L. Y. A., Lancaster, L., et al. 2025, arXiv e-prints, arXiv:2505.05442
 - Stanway & Eldridge (2018) Stanway, E. R. & Eldridge, J. J. 2018, MNRAS, 479, 75
-- Tacchella et al. (2023) Tacchella, S., Eisenstein, D. J., Hainline, K., et al. 2023, ApJ, 952, 74
+- Strait et al. (2023) Strait, V., Brammer, G., Muzzin, A., et al. 2023, ApJ, 949, L23
 - Treu et al. (2022) Treu, T., Roberts-Borsani, G., Bradac, M., et al. 2022, ApJ, 935, 110
@@
 - Vulcani et al. (2017) Vulcani, B., Trenti, M., Calvi, V., et al. 2017, ApJ, 836, 239
-- Weibel et al. (2024) Weibel, A., de Graaff, A., Setton, D. J., et al. 2024, arXiv e-prints, arXiv:2409.03829
-- Westcott et al. (2024) Westcott, L., Conselice, C. J., Harvey, T., et al. 2024, arXiv e-prints, arXiv:2412.14970
+- Weibel et al. (2025a) Weibel, A., de Graaff, A., Setton, D. J., et al. 2025a, ApJ, 983, 11
+- Weibel et al. (2025b) Weibel, A., Oesch, P. A., Williams, C. C., et al. 2025b, arXiv e-prints, arXiv:2507.06292
+- Westcott et al. (2025) Westcott, L., Conselice, C. J., Harvey, T., et al. 2025, ApJ, 983, 121
 - Whitler et al. (2025) Whitler, L., Stark, D. P., Topping, M. W., et al. 2025, arXiv e-prints, arXiv:2501.00984
@@
 We present here the extended samples of z (ID, coordinates and F356W magnitudes from M24 in Table 4, SED, thumbnails and in Fig.
-13 and 14) and z (Table 5, Fig.
-15 and 16) candidates.
+12 and 13) and z (Table 5, Fig.
+14 and 15) candidates.
 We remark that while these objects also have colours and consistent with high-redshift solutions, and shall be considered of interest for potential spectroscopic follow-up, the contamination level in these samples is expected to be significant (Sect.
@@
 | UDS_130776 | 34.434420 | -5.108762 | 28.35 0.14 |
+Appendix B Spectral energy distributions of the confirmed interlopers
+Main properties (Table 6) and SEDs (Fig.
+16) of the six known interlopers of the F200W-dropout selection.
+Objects COSMOS_31168, COSMOS_35731, COSMOS_76919 and UDS_56824 have been found to be at z2-7 on the basis of NIRSpec observations carried out by the CAPERS survey (Sec.
+4.3).
+Object A2744_27713 is a transient which is detectable in the second-epoch GLASS-JWST observations (November 2022) when its position was covered only by NIRCam LW bands.
+| ID | R.A. | Dec | zspec | F356W | zhigh | zlow | Type | ||
+| deg. | deg. | AB | |||||||
+| A2744_27713 | 3.522866 | -30.351297 | - | 28.56 0.07 | 18.7 | 2.9 | 0.46 | 3.28 | 1 |
+| CEERS_15937 | 214.944272 | 52.835847 | 4.91 | 27.15 0.09 | 17.2 | 4.6 | 0.27 | 0.84 | 2 |
+| CEERS_81406b | 214.914542 | 52.943021 | 4.91 | 26.52 0.03 | 16.3 | 4.6 | 0.94 | 3.90 | 3 |
+| COSMOS_31168 | 150.180936 | 2.260756 | 2.29 | 27.67 0.16 | 16.2 | 4.0 | 2.85 | 1.31 | 3 |
+| COSMOS_35731 | 150.133571 | 2.271020 | 3.66 | 28.54 0.09 | 15.9 | 2.8 | 1.07 | 0.88 | 3 |
+| COSMOS_76919 | 150.184553 | 2.353510 | 5.18 | 28.71 0.18 | 16.1 | 4.0 | 1.76 | 0.74 | 3 |
+| UDS_56824 | 34.454893 | -5.215586 | 6.56 | 26.51 0.12 | 19.8 | 4.2 | 1.45 | 1.93 | 3 |
+a) ID, coordinates and F356W magnitudes from M24.
+The best-fit photometric redshift solutions have been obtained with zphot at () and ().
+The last column report the classification discussed in Sect.
+3.2 and 4.2 (1=transient; 2=quiescent; 3=dusty star-forming).
+b) Object CEERS-93316 from Arrabal Haro et al. (2023b).